### v1

# Large Language Models as Reliable Knowledge Bases?

Abstract
The NLP community has recently shown a growing interest in leveraging Large Language Models (LLMs) for knowledge-intensive tasks, viewing LLMs as potential knowledge bases (KBs). However, the reliability and extent to which LLMs can function as KBs remain under-explored. While previous studies suggest LLMs can encode knowledge within their parameters, the amount of parametric knowledge alone is not sufficient to evaluate their effectiveness as KBs. This study defines criteria that a reliable LLM-as-KB should meet, focusing on factuality and consistency, and covering both seen and unseen knowledge. [^1] Seen knowledge refers to knowledge learned during training. Unseen knowledge is neither present in the model’s training data nor can be inferred from seen knowledge. We develop several metrics based on these criteria and use them to evaluate 26 popular LLMs, while providing a comprehensive analysis of the effects of model size, instruction tuning, and in-context learning (ICL). Our results paint a worrying picture. Even a high-performant model like GPT-3.5-turbo is not factual or consistent, and strategies like ICL and fine-tuning are unsuccessful at making LLMs better KBs.
1 Introduction
Large language models (LLMs), pretrained on extensive text corpora, have demonstrated the ability to implicitly encode various types of knowledge within their weights, without requiring human supervision. As a result, many recent studies Chuang et al. (2023); Yu et al. (2023); Dhingra et al. (2022); Sung et al. (2021); Wang et al. (2020) aim to analyze the relationship between LLMs and KBs, and even explore whether LLMs can replace KBs Sun et al. (2023); Mruthyunjaya et al. (2023); Heinzerling and Inui (2021).
However, whether current LLMs can serve as reliable KBs and how to evaluate their performance in this role remains largely unexplored. Existing studies Sun et al. (2023); Wang et al. (2021); Roberts et al. (2020) often implicitly assume that the LLM’s ability to retain knowledge is sufficient for it to function as a KB. Typically, these studies employ two methods: (1) converting knowledge graphs into natural language questions using templates and evaluating LLM ability to answer these questions, by measuring the amount of knowledge therein Petroni et al. (2019); Sun et al. (2023); and (2) pre-training LLMs on passages/triples containing knowledge and then assessing their ability to answer related questions, thereby quantifying their knowledge retention Wang et al. (2021); He et al. (2024). While these methods demonstrate that LLMs can recall knowledge, memorizing facts is not the sole criterion for being a reliable KB AlKhamissi et al. (2022).
What criteria then should a LLM meet to function as a reliable KB? Discussion on this topic has been limited, and there is no agreement on the definition of these criteria. AlKhamissi et al. (2022) argue that LLMs ought to excel at five aspects (i.e., access, edit, consistency, reasoning, explainability, and interpretability) if they are to be considered a KB. However, they do not outline specific metrics to evaluate the extent to which LLMs act as KBs. Besides, we argue that evaluating LLMs against the characteristics of KBs may not be entirely appropriate due to their different data storage structures. Instead, we should consider the specific properties of LLMs when assessing their suitability as KBs.
Our research addresses these gaps and aims to establish a more nuanced understanding of LLMs-as-KBs. We evaluate and compare the reliability of different LLMs functioning as KBs in answering factoid questions. Specifically, our work makes the following contributions:
  1. 1.
    We define how to assess the reliability of LLMs-as-KBs, and propose metrics that consider the differences between KBs and LLMs and align with the distinct nature of LLMs. We consider two aspects, factuality (i.e., the ability to provide factual responses) and consistency (i.e., the ability to provide consistent responses for questions involving the same knowledge), and cover the evaluation on both seen and unseen knowledge.
  2. 2.
    To evaluate consistency, we propose a novel method which computes the probability that a LLM can consistently provide response $r$ for question $q$.
  3. 3.
    To evaluate LLMs on unseen knowledge, we create a new QA dataset UnseenQA, with a knowledge cutoff date before April 13, 2024.
  4. 4.
    We evaluate 26 popular LLMs in their ability to function as reliable KBs, and discuss the influence of model size, instruction-tuning, and ICL on their performance.
Figure 1: An example illustrating three answer types: correct, wrong, and uninformative. Focusing only on the correct rate incorrectly suggests that Model B is better, even though Model A is more reliable with a similar correct rate and a much lower wrong rate.
2 What is a Reliable LLM-as-KB?
In simple terms, a LLM functions as a reliable KB if it consistently provides factual responses. Evaluating the reliability of LLMs as KBs primarily involves assessing two critical dimensions, namely factuality and consistency.
2.1 Factuality
Factuality refers to the quality of being factual or based on fact. Run-of-the-mill KBs, stored on physical servers or cloud platforms, deliver information directly in response to queries. If the requested data is unavailable, these systems typically return a null response. In contrast, LLMs are probabilistic models that excel at next word prediction based on the given context, rather than storing explicit information in defined locations. This architecture allows LLMs to generate responses that seem plausible, regardless of whether the content was included in their training data. Consequently, LLMs typically produce three types of responses: correct, uninformative, and wrong.
Ongoing efforts Chen et al. (2023); Wang et al. (2024) to evaluate the factuality of LLMs often hinge on the models’ correct rates in responding to factual QA datasets. However, this approach has notable limitations. Firstly, many studies Lin et al. (2022); Sun et al. (2023) do not specify whether the dataset’s scope of knowledge was included in the LLM’s training data. This omission can lead to unfair comparisons between models, especially if the knowledge being tested is within the training scope of one model but not another. Secondly, the assumption that a higher correct rate indicates greater factuality is problematic. For example, consider models A and B shown in Figure 1, and assume they are being evaluated on a test dataset covering knowledge they both have seen during training. In this case, focusing solely on the correct rate might erroneously suggest that Model B is more factual, despite Model A being more reliable due to its similar correct rate and significantly lower wrong rate.
Given these issues, we propose the following criteria for the factuality of LLMs-as-KBs:
CRITERION 1.1
For seen knowledge, a factual LLM should demonstrate a high correct rate and a low wrong rate.
CRITERION 1.2
For unseen knowledge, a factual LLM should demonstrate a high uninformative rate.
We next proceed to define evaluation metrics that operationalize these criteria. Let $\boldsymbol{M}$ denote a LLM. Let $D_{\text{seen}}$ denote a QA dataset containing $N$ open-ended factoid questions pertaining to knowledge the LLM ought to have seen during training. Let $D_{\text{unseen}}$ denote a QA dataset with $L$ open-ended factoid questions covering unseen knowledge. We further assume the LLM’s response to $D_{\text{seen}}$ will be correct, uninformative, or wrong, while its response to $D_{\text{unseen}}$ will be either uninformative or wrong.
METRIC 1.1: Net Correct Rate (NCR)
measures how much more likely the model is to provide correct responses instead of wrong ones on $D_{\text{seen}}$ questions. It is defined as:
[…]
$\text{CR}=\frac{N_{correct}}{N}$ $\text{WR}=\frac{N_{wrong}}{N}$ (2)
where $N_{correct}$ and $N_{wrong}$ are counts of correct and wrong responses, respectively.
NCR values range from $-1$ to $1$. A negative NCR suggests the model tends to provide misleading responses, while a positive NCR suggests a preference for correct responses. Consider again two models, A and B. According to CRITERION 1.1, if model A has a higher correct rate and lower wrong rate compared to model B, then model A is better. Formally, if $\text{CR}_{A}-\text{CR}_{B}>\text{WR}_{A}-\text{WR}_{B}$, then model A is better than B. Algebraically, this is equivalent to $\text{CR}_{A}-\text{WR}_{A}>\text{CR}_{B}-\text{WR}_{B}$, i.e., $\text{NCR}_{A}>\text{NCR}_{B}$. Therefore, a higher NCR indicates a more factual model on seen knowledge.
METRIC 1.2: Uninformative Rate (UR)
assesses whether the model is likely to provide uninformative responses to $D_{\text{unseen}}$ questions. It is formulated as:
$\text{UR}=\frac{L_{uninformative}}{L}$ (3)
where $L_{uninformative}$ denotes the count of uninformative responses. UR ranges from 0 to 1. A higher UR indicates that the model is more likely to refrain from giving wrong responses when faced with unseen knowledge.
Figure 2: Illustration of LLM inconsistency with davinci-002 (temperature is set to 0). Questions in the top focus on seen knowledge, with probability distribution mass concentrated on one prediction. Questions in the bottom focus on unseen knowledge, where the distribution is more even. Drawing from such a distribution inevitably leads to inconsistencies.
2.2 Consistency
Consistency refers to the quality of always behaving in the same way or having the same opinions. KBs are designed with this principle in mind. In fact, there are efficient algorithms Andersen and Pretolani (2001) which detect and resolve conflicts within KBs, thus ensuring consistent outputs for queries on the same facts.
It is well-known that LLMs often exhibit inconsistencies in their responses Elazar et al. (2021); Wang et al. (2022). Current research Elazar et al. (2021); Jang et al. (2022); Hagström et al. (2023) evaluates LLM consistency through their performance on benchmarks involving paraphrasing, negation, or multilingual variations. A model is considered superior if it responds consistently across a broader range of data samples. In this work, we argue that it may be too strict to expect LLMs to be always consistent when responding to fact-based questions. Unlike explicit KBs which store information at a fixed location, LLMs operate probabilistically. In theory, if the context has been learned during training, the probability distribution for the prediction will be concentrated; otherwise, it will be more uniform. Drawing from a uniform probability distribution inevitably leads to inconsistencies. As shown in Figure 2, even with greedy decoding, slight biases in the distribution can cause fluctuations in the selection of the top probable words.
Given their probabilistic nature, we do not expect LLMs to always behave consistently. We acknowledge that inconsistencies can cause confusion in practical applications and propose to monitor model behavior through post-processing which we argue is more realistic than expecting a probabilistic model to be perpetually consistent. We thus propose the following consistency criteria:
CRITERION 2.1
The model is expected to be consistent when it produces correct responses.
CRITERION 2.2
The model is expected to be inconsistent when it produces wrong responses.
We next define evaluation metrics corresponding to the criteria above. Let $q$ refer to a question in either $D_{\text{seen}}$ or $D_{\text{unseen}}$, and $r$ denote model $M$’s response to $q$. Inspired by Zheng et al. (2024), we measure consistency based on multiple-choice questions (MCQs). As shown in Figure 3, we employ gpt-3.5-turbo-insruct to generate a set of distractor options similar to response $r$, and then create a group of MCQs. The consistency score for data point $(q,r)$ is calculated as:
$Cons(q,r)=\frac{\sum_{i=1}^{X_{\text{MCQs}}}[R_{i}=r]}{X_{\text{MCQs}}}$ (4)
where $X_{\text{MCQs}}$ is the total number of MCQs, $R_{i}$ is model $M$’s response for the $i$-th MCQ, and the expression $[R_{i}=r]$ yields 1 when the model’s response $R_{i}$ matches its original response $r$, and 0 otherwise. The consistency score $Cons(q,r)$ ranges from 0 to 1.
METRIC 2.1: $C_{\textit{correct}}$
measures the consistency of the model when it provides correct responses. It is defined as:
$C_{\textit{correct}}=\frac{\sum_{j=1}^{N_{correct}}Cons(q_{j}^{(c)},r_{j}^{(c)})}{N_{correct}}$ (5)
where $r^{(c)}$ refers to the response labeled as correct, and $q^{(c)}$ is the corresponding question. $C_{\textit{correct}}$ ranges from 0 to 1. Based on CRITERION 2.1, a higher $C_{\textit{correct}}$ is desirable.
METRIC 2.2: $C_{\textit{wrong}}$
measures the consistency of an LLM when it provides wrong responses. It is defined as:
$C_{\textit{wrong}}=\frac{C_{\textit{wrong}}^{s}+C_{\textit{wrong}}^{u}}{2}$ (6)
where $C_{\textit{wrong}}^{s}$ refers to the consistency of a LLM when it provides wrong responses to questions about seen knowledge, and $C_{\textit{wrong}}^{u}$ refers to the consistency of a LLM when it provides wrong responses to questions about unseen knowledge:
$C_{\textit{wrong}}^{s}=\frac{\sum_{j=1}^{N_{wrong}}Cons(q_{j}^{(w)},r_{j}^{(w)})}{N_{wrong}}$ (7)
$C_{\textit{wrong}}^{u}=\frac{\sum_{j=1}^{L_{wrong}}Cons(q_{j}^{(w)},r_{j}^{(w)})}{L_{wrong}}$ (8)
where $r^{(w)}$ refers to responses labeled as wrong, and $q^{(w)}$ is the corresponding question. $N_{wrong}$ and $L_{wrong}$ are counts of wrong responses to $D_{\text{seen}}$ and $D_{\text{unseen}}$, respectively. $C_{\textit{wrong}}$ ranges from 0 to 1. Based on CRITERION 2.2, a lower $C_{\textit{wrong}}$ is desirable.
Figure 3: Example computation for consistency score $Cons(q,r)$. The LLM’s original answer is shown in green, while distractors are red.
2.3 Reliability (Factuality and Consistency)
A reliable LLM-as-KB should then be assessed against factuality and consistency. Based on the criteria defined above, a LLM-as-KB is reliable if it meets the following:
CRITERION 3.1
For seen knowledge, a LLM should have a high rate of consistently correct responses and a low rate of consistently wrong responses.
CRITERION 3.2
For unseen knowledge, a LLM should have a high rate of uninformative or inconsistent responses.
We quantify these criteria with two metrics.
METRIC 3.1: Net Consistently Correct Rate (NCCR)
quantifies the model’s tendency to provide consistently correct responses compared to consistently wrong ones for questions about seen knowledge. It is defined as:
[…]
CCR $=\text{CR}\times C_{correct}$
CWR $=\text{WR}\times C_{wrong}^{s}$
NCCR ranges from $-1$ to $1$. A higher NCCR indicates a LLM is more reliable on seen knowledge. A negative NCCR suggests the model provides consistently wrong responses, while a positive NCCR suggests a preference for consistently correct responses.
METRIC 3.2: Inconsistent/Uninformative Rate (IUR)
assesses whether the model is likely to provide uninformative or inconsistent wrong responses for questions about unseen knowledge. It is defined as:
$\text{IUR}=1-(1-\text{UR})C_{wrong}^{u}$ (10)
IUR ranges from 0 to 1. A higher IUR value indicates the LLM functions as a more reliable KB on unseen knowledge.
LLM Zero-Shot Four-Shot Four-Two
NCCR IUR AVG NCCR IUR AVG NCCR IUR AVG
gpt-3.5-turbo 34.1 95.7 62.7 35.5 99.0 66.7 32.1 99.8 65.8
llama2chat-70b 20.4 98.9 59.1 17.7 99.9 58.8 16.4 100.0 58.2
llama3-70b 30.2 71.6 36.7 33.1 69.2 35.7 33.9 98.8 65.8
Table 1: Most reliable LLMs across three prompt settings. Four-Two refers to the four-shot with two unsure shots setting. All numbers shown are percentages. AVG represents average NCCR and (normalized) IUR ($2*\text{IUR}-100$) scores (both scores use the same scale). We treat seen and unseen knowledge equally, however, in practice, NCCR and IUR can be assigned different weights.
3 Experimental Setup
3.1 Datasets
To evaluate the performance of LLMs on seen knowledge, we collated SeenQA, a composite dataset containing 3,000 questions sourced from the test sets (or development sets, where test sets were unavailable) of Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), and PopQA Mallen et al. (2023). For assessing LLM performance on unseen knowledge, we introduce UnseenQA, a new dataset designed to ensure that models with a knowledge cutoff date before April 13, 2024 do not have access to answers. UnseenQA also includes 3,000 questions, which are derived from 20 templates covering various answer types such as number, people, time, location, and others. Our methodology for creating SeenQA and UnseenQA is detailed in Appendix A.
3.2 LLM Selection
We evaluate a wide range of popular LLMs (26 in total) in their ability to function as KBs and investigate the impact of various factors, such as the number of parameters and fine-tuning. Specifically, we consider the following models: gpt-3.5-turbo, flan-t5 (0.08B, 0.25B, 0.78B, 3B, 11B), llama1 (7B, 13B, 65B), llama2 (7B, 13B, 70B), llama2-chat (7B, 13B, 70B), llama3 (8B, 70B), llama3instruct (8B, 70B), mistral (7B), mistral-instruct (7B), gemma (2B, 7B), gemma-instruct (2B, 7B), and Phi2 (3B). Detailed descriptions of these models are provided in Table 4 in Appendix B. We classify LLMs into three categories based on their parameter sizes: small (0.08B–3B), medium (7B–13B), and large (65B–70B). We use the term ‘fine-tuned LLMs’ to refer to the LLMs that have been fine-tuned through instruction-tuning or reinforcement learning from human feedback. We use the term ‘base LLMs’ to refer to LLMs without fine-tuning.
3.3 Evaluation on a Single Response
Uninformative
We identify three types of uninformative responses from LLMs: ‘repetition’, ‘none’, and ‘unsure’. ‘Repetition’ refers to responses that repeatedly echo a specific string. We detect this using regular expressions and word frequency analysis. ‘None’ denotes responses lacking relevant information, such as an empty string and repetition of the question. ‘Unsure’ indicates responses where the model explicitly states it is unable to answer or lacks the required knowledge. We label a response as ‘unsure’ if it includes expressions such as ‘I am not sure’, ‘I cannot provide’, and ‘I am just an AI’.
Correct
We determine whether a response is correct based on exact match. A response is considered correct if the exact match score is 1. In cases where exact match is 0, we compare the model’s prediction against the ground truth using gpt-4o. Previous work Sun et al. (2023) has shown a 98% agreement rate between ChatGPT and human judgments in comparing model responses with ground truth. We follow their prompt as outlined in Table 7 in Appendix C.
Consistency Score
To compute $Cons(q,r)$, we set $X_{\text{MCQs}}$ (total number of MCQs) to 20, and each MCQ includes question $q$ and 5 options (the original response $r$, 3 random distractor options, and an ‘unsure’ option).
3.4 Prompts and Hyper-parameters
To provide a comprehensive evaluation, we experimented with three types of prompt settings: zero-shot, four-shot, and four-shot with two unsure shots. To avoid any bias introduced by fixed examples, we employed a dynamic few-shot method following the work of Nori et al. (2023). We collected two repositories, $R_{seen}$ and $R_{unseen}$. $R_{seen}$ includes 280 question-answer pairs about seen knowledge (200 from the unused data of PopQA and training data of Natural Questions and TriviaQA; 80 are generated using the templates in Table 3). $R_{unseen}$ consists of 40 question-answer pairs about unseen knowledge, all generated using the templates in Table 3. We used text-embedding-3-small to embed the questions in the repositories and test questions as vector representations. For each test question under the four-shot setting, we retrieved its nearest four questions from $R_{seen}$. Under the four-shot with two unsure shots setting, we retrieved the nearest two questions from $R_{seen}$ and two from $R_{unseen}$.
All LLMs were evaluated using greedy decoding (temperature 0 for gpt-3.5-turbo) with a maximum of 100 new tokens. Our prompts are provided in Appendix C.
Figure 4: Comparing llama3instruct-8b against llama2chat-13b, and phi2-3b against gemma-2b using CR, WR, and NCR metrics (zero-shot setting). Even though these models are comparable according to CR, they are quite different according to WR and NCR metrics. We observe similar trends in the four-shot and with two unsure shots settings.
4 Results
We present detailed results for all LLMs are in Table 9 (factuality), Table 10 (consistency), and Table 11 (reliability) in Appendix D. LLM rankings based on different metrics are shown in Figure 12 and Figure 13, also in Appendix D.
gpt-3.5-turbo is overall the most reliable LLM.
Table 1 presents results for the two best performing LLMs under different prompt settings. As can be seen, gpt-3.5-turbo is most reliable across the board. Although it is not consistently wrong when asked about facts it does not know (its IUR score exceeds 95%), it is not consistently correct when asked about facts it has seen before (NCCR is only 32%).
Figure 5: LLM ranking based on NCCR and IUR metrics in zero-shot setting. See full results in Figure 12 (Appendix D) for other shot settings.
flan-t5-0.78B is the most reliable LLM for unseen knowledge and most unreliable with seen knowledge.
Figure 5 shows the LLMs ranked by NCCR and IUR in a zero-shot setting. As can be seen, while flan-t5-0.78B ranks low for NCCR, it maintains the top position for IUR. A similar trend is observed with the flan-t5-3B and gemma-instruct (2B, 7B) models. Conversely, models in the llama3 family show the opposite trend: they rank high on seen knowledge but low on unseen knowledge. For instance, llama3-70b, despite ranking second in terms of NCCR, falls to the fourth lowest position for IUR.
Net Correct Rate (NCR) reveals factuality gaps in LLMs.
As illustrated in Figure 4, CR, the standard metric for assessing factuality, fails to fully capture nuanced differences. Despite similar CR values of approximately 40%, llama3instruct-8B and llama2chat-13B behave differently when it comes to wrong responses (on seen data); the former model has a WR of 15% higher than the latter, and as a result its NCR is substantially lower. In the case of phi2-3B, Gunasekar et al. (2023) claim that with "textbook quality" data, smaller LLMs can achieve satisfactory performance with less training data. Focusing solely on CR, this claim is true as phi2-3B’s CR is comparable to that of gemma-2B, which was trained with twice the token count. However, phi2-3B exhibits significantly higher WR (and lower NCR) compared to gemma-2B which underscores the challenge of maintaining low error rates even with superior data quality.
A less factual LLM can be more reliable.
Figure 6 illustrates this finding. The llama3instruct-8b model has a slightly worse NCR compared to llama2chat-7B, indicating that the former model is marginally less factual on seen knowledge. However, llama3instruct-8b exhibits a higher NCCR due to its better consistency score on correct responses. This suggests that llama3instruct-8b is more reliable than llama2chat-7B on seen knowledge. Additionally, both LLMs show negative NCR and positive NCCR in the zero-shot setting. This indicates that although they produce more wrong than correct responses, they generate fewer consistently wrong responses compared to consistently correct ones.
Figure 6: Comparison between llama3instruct-8b and llama2chat-7b (zero-shot setting). The former model is slightly less factual according to NCR, but more consistent in its responses (see NCCR, $C_{correct}$, and $C^{s}_{wrong}$). We observe similar trends in the four-shot and with two unsure shots settings.
Comparisons Zero-Shot Four-Shot Four-Two
NCR vs. UR 0.27 0.34 0.62
NCCR vs. IUR --0.17 --0.12 0.41
$C_{\textit{correct}}$ vs. $C_{\textit{wrong}}$  0.81  0.78 0.51
Table 2: Pairwise correlation of LLM performance on different metrics under different prompt settings. The correlation values are computed across all LLMs (26 data points). We report Pearson’s $\rho$, diacritic  denotes statistical significance ($p<0.05$). Four-Two refers to the four-shot setting with two unsure shots.
Performance on seen knowledge is not predictive of performance on unseen knowledge.
In Table 2, we examine whether metrics applied to seen knowledge can be used to extrapolate model performance on unseen knowledge by reporting correlation values (Pearson’s $\rho$) between NCR and UR, and NCCR and IUR. As can be seen, in both zero- and four-shot settings, correlations are not statistically significant. Nevertheless, correlations are significant in all metric comparisons in the four-shot with two unsure shots setting. While in general model performance with seen knowledge does not transfer to unseen knowledge, specific prompt manipulations can enhance the correlation between metrics (see last column in Table 2).
LLMs are consistently right and wrong!
Table 2 reports a positive, significant correlation between $C_{correct}$ and $C_{wrong}$. This result implies that LLMs demonstrating high consistency in correct responses also tend to exhibit high consistency in wrong responses. This finding contradicts our expectation of high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$ and highlights a notable flaw in current models. Future work should address this issue, e.g., by instructing LLMs to achieve the desired consistency behavior.
Figure 7: The impact of model size on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics are color-coded. LLMs are shown in three sizes, small, medium, and large and are grouped into ‘base’ and fine-tuned ones.
5 Analysis
In this section, we explore the effects of model size, fine-tuning, and ICL on LLM performance based on different metrics. We also analyze how LLMs handle unseen knowledge by looking at the distribution of uninformative responses and the impact of different question types.
Figure 8: The impact of fine-tuning on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics color-coded. This analysis only considers the performance of Llama2, Llama3, Mistral, and Gemma as these families include both base LLMs and fine-tuned versions. Models are shown in three sizes, small, medium, and large.
Figure 9: The impact of ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR. Different metrics are color-coded. We compare zero-shot and four-shot settings; and zero-shot against four-shot with two unsure shots. LLMs (‘base’ and fine-tuned ones) are in three sizes, small, medium, and large.
Figure 10: The impact of question type on LLM performance as measured by Uninformative Rate (UR) on unseen knowledge. Questions are grouped into types based on answer they require. Higher values have darker shades.
5.1 The Impact of Model Size
Larger LLMs perfom better on seen knowledge but worse on unseen knowledge.
Figure 7 demonstrates that as model size increases, both NCR (blue line) and NCCR (purple line) improve. This trend indicates that larger LLMs perform better on questions about seen knowledge. However, for base LLMs in both zero-shot and four-shot settings, IUR (brown line) decreases as model size increases. This trend suggests that as LLMs become larger, they become more consistent at delivering wrong responses for questions regarding unseen knowledge.
Larger LLMs are more consistent, even with wrong responses.
In Figure 7, we observe that as model size increases, both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) increase significantly. While higher consistency scores for correct responses are expected and useful, the higher consistency scores for wrong responses pose a potential risk. Larger models may consistently produce convincing but wrong information, which could lead to misinformation if not carefully managed.
5.2 The Impact of Fine-tuning
Fine-tuning improves performance on unseen knowledge but negatively affects performance on seen knowledge.
As illustrated in Figure 8, UR (orange line) and IUR (brown line) show significant improvement in LLMs after fine-tuning, indicating an enhanced ability to distinguish and respond to unseen knowledge appropriately. However, the decreasing NCCR (purple line) suggests they become worse at handling seen knowledge after fine-tuning.
Fine-tuning does not make LLMs more consistent.
As depicted in Figure 8, there is no noticeable increase in $C_{\textit{correct}}$ (green line) after fine-tuning. In fact, fine-tuning even has a negative impact on $C_{\textit{correct}}$ for medium-sized LLMs. These results indicate that current fine-tuning techniques fail to enhance the consistency of correct responses. Furthermore, $C_{\textit{wrong}}$ (red line) does not decrease after fine-tuning either, which suggests that fine-tuning also fails to reduce the model’s presistence on wrong responses.
5.3 The Impact of ICL
Unsure shots improve LLM performance on unseen knowledge.
As shown in Figure 9, incporating two unsure shots in the four-shot setting substantially increases UR (orange line) and IUR (brown line) across model sizes. In contrast, performance on unseen knowledge deteriorates when using four-shots only for all sizes of base LLMs and small fine-tuned LLMs, compared to their zero-shot counterparts.
ICL does not improve LLM perfomance on seen knowledge.
Figure 9 demonstrates that ICL does not improve LLM performance according to NCR (blue line) or NCCR (purple line) metrics. For small/medium base LLMs and small fine-tuned LLMs in the four-shot setting, ICL even decreases NCR performance. This finding contrasts with previous results about the ability of LLMs to learn from a few examples Brown et al. (2020); Chada and Natarajan (2021); Touvron et al. (2023); Bai et al. (2023). Earlier work has mostly focused on the correct rate of LLMs in few-shot settings without unsure shots. Based on the results presented in Table 9 in Appendix D, we observe that for most LLMs, the four-shot setting indeed substantially improves the correct rate but also increases the wrong rate. Incorporating unsure shots helps decrease the wrong rate but also decreases the correct rate.
ICL makes LLMs less consistent, even on correct responses!
Figure 9 shows that ICL leads to reductions in both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) for medium/large fine-tuned LLMs and large base LLMs. While a lower $C_{\textit{wrong}}$ is desirable, the decrease in $C_{\textit{correct}}$ is concerning. This indicates that ICL negatively affects the reliability of LLMs at providing correct responses.
Figure 11: Distribution of uninformative responses given by LLMs to questions about unseen knowledge. We report results for the llama3-8b, gemma-7b, and their fine-tuned models (fourth column) but observe similar trends on other models (omitted for the sake of brevity).
5.4 Model Behavior on Unseen Knowledge
Base LLMs overestimate their knowledge on numerical and temporal questions.
Figure 10 reports UR performance across models, broken down per question type (e.g., number, person, location). As can be seen, most base LLMs exhibit significantly lower UR performance on questions requiring numerical or temporal responses. This indicates that base LLMs often provide misleading information for such questions, even when they lack the relevant knowledge.
Fine-tuned LLMs can explicitly admit when they don’t know.
As illustrated in Figure 11 (fourth column), fine-tuned LLMs are able to explicitly acknowledge their lack of knowledge by answering ‘unsure’ (see definition in Section 3.3) to questions about unseen knowledge. In contrast, base LLMs often produce responses classified as ‘none’ or ‘repetition’ in the absence of unsure shots (contrast columns one and two with column three in Figure 11).
6 Related Work
Petroni et al. (2019) were the first to introduce the concept of using pre-trained language models as knowledge bases. They introduced LAMA, a benchmark with questions structured as "fill-in-the-blank" cloze statements, and found that BERT Devlin et al. (2019) retains relational knowledge competitive with traditional NLP methods that have some access to oracle knowledge.
Roberts et al. (2020) were the first to measure the extent to which language models trained on unstructured text can implicitly store and retrieve knowledge using natural language queries. Specifically, they fine-tuned T5 Raffel et al. (2019) to answer questions without access to any external context or knowledge and showed that this approach performs competitively with open-domain systems that explicitly retrieve answers from an external knowledge source when answering questions. On a similar vein, Wang et al. (2021) fine-tuned BART Lewis et al. (2020) with related passages to instill factual knowledge and used the accuracy of masked span recovery to measure the extent to which this knowledge was memorized by the model. They found it was challenging for pre-trained language models like BART to remember facts seen in training, and answer questions, even in cases where the relevant knowledge was retained. He et al. (2024) explicitly trained T5 Raffel et al. (2019) and LLAMA2 Touvron et al. (2023) to memorize world knowledge from Wikidata on a large scale and then used exact match and F1 scores to evaluate knowledge retention. Their results showed LLMs hold promise as large-scale KBs capable of retrieving facts and responding with flexibility, but are less proficient at inferring new knowledge through reasoning.
Sun et al. (2023) proposed a benchmark consisting of 18,000 question-answer pairs representing facts with varying popularity (i.e., high, medium, low) and assessed the knowledge retained by various language models on this benchmark. Unlike previous studies, Sun et al. (2023) reported both accuracy (the percentage of questions answered correctly) and hallucination rate (the percentage of questions answered wrongly). They showed models are particularly bad at answering questions from medium- and low-frequency facts.
Our work also examines whether LLMs can be used in lieu of more traditional KBs. Compared to previous research, we proposed a comprehensive evaluation framework which assesses not only the ability of LLMs to recall seen knowledge but also their ability to respond in the face of unseen knowledge. In addition, we evaluate LLM consistency when answering questions about identical knowledge.
7 Conclusion
In this paper, we defined a set of criteria that LLMs functioning as KBs should meet, focusing on factuality and consistency. We proposed various metrics operationalizing these criteria and used them to assess LLM performance when answering questions pertaining to both seen and unseen knowledge. We evaluated 26 popular LLMs and found that gpt-3.5-turbo is the most reliable among them. Additionally, we examined the impact of model size, fine-tuning, and ICL. Fine-tuning and ICL with unsure shots were shown to improve LLM capabilites when providing responses to questions about unseen knowledge, but do not significantly improve performance on seen knowledge. Increasing model size boosts performance on seen knowledge but at the expense of performance on unseen knowledge. Notably, neither ICL nor fine-tuning were successful at improving LLM consistency. This highlights the critical need for continued research to develop more robust strategies that ensure both factuality and consistency, enabling LLMs to effectively function as KBs.
Appendix A Datasets
A.1 SeenQA Dataset
SeenQA is composed of questions selected from the following datasets:
  1. 1.
    Natural Questions Kwiatkowski et al. (2019): This dataset includes questions sourced from web queries, each paired with a corresponding Wikipedia article containing the answer. The paper on Natural Questions was submitted to TACL in April 2018.
  2. 2.
    TriviaQA Joshi et al. (2017): This dataset comprises questions from Quiz League websites, supplemented by web pages and Wikipedia searches that may contain the answer. The paper on TriviaQA was submitted to Arxiv in May 2017. For this project, we focus only on questions supported by Wikipedia.
  3. 3.
    PopQA Mallen et al. (2023): This dataset targets long-tail entities. The authors used the Wikipedia dump from December 2018 in the retrieval augmented baseline, indicating that the knowledge in PopQA can be covered by the Wikipedia dump from that date.
Wikipedia is a common source in the pre-training data of large language models (LLMs). Comparing the knowledge cutoff dates provided in Table 4, we can deduce that the knowledge involved in these three datasets must have been seen during training by the LLMs used in our study. SeenQA contains 3000 questions selected through a three-step process:
  1. 1.
    Factoid Question Extraction: We filter out questions starting with "why," those with multiple different answers, or answers longer than five tokens.
  2. 2.
    Removal of Time-Sensitive Questions: We use gpt-4-1106-preview to detect questions whose answers may change over time, using the prompt shown in Table 5. Such questions are then removed from the dataset.
  3. 3.
    Sampling: Finally, we randomly sample 1,000 questions from each dataset’s test or development set, if the test set is unavailable, discarding the supporting context to adapt to a closed-book setting.
A.2 UnseenQA Dataset
UnseenQA was created using 20 templates, shown in Table 3, covering five answer types: number, person, time, location, and others. These templates wet=re hand-written by the authors. Templates T1, T2, T5, T6, and T7 are about future events, so their answers are unknown at the time of writing. The remaining templates involve fictional persons whose names and related information cannot be found on the Internet. The templates include three types of placeholders:
  1. 1.
    Country/Region: We collected 150 country/region names from the list of National Olympic Committees on a Wikipedia page.
  2. 2.
    Medal Event: We collected 150 medal events from the official programme of the Olympic games, Paris 2024.
  3. 3.
    Person: We randomly sampled 150 person names from combinations of 100 first names, 100 middle names, and 100 last names. We manually checked these names on Google to ensure they have no Internet presence.
UnseenQA consists of 3000 questions, with 150 questions per template. This dataset was created on April 13, 2024. Therefore, LLMs trained on knowledge before this date should not know the answers.
Answer Type Abb Template
Number T1 How many gold medals did [country/region] win at the XXXIV Summer Olympic Games?
T2 In the 25th FIFA World Cup, what was the final ranking of [country/region]?
T3 How many children does [person] have?
T4 How old was [person] in 2015?
Person T5 Who won the bronze medal of [medal event] at the XXXIII Summer Olympic Games?
T6 Who is the supreme leader of [country/region] in 2040?
T7 In 2028, who served as the head coach of [country/region] national football team?
T8 Who is [person]’s mom?
Time T9 On which date was [person] born?
T10 In what year did [person] die?
T11 In what year did [person] graduate with the bachelor’s degree?
T12 When was the wedding date for [person]?
Location T13 Where was [person] born?
T14 Where did [person] pass away?
T15 Which university did [person] attend for the undergraduate studies?
T16 Where was [person]’s wedding held?
Others T17 What was the cause of [person]’s death?
T18 What is the title of the debut album released by [person]?
T19 What is the name of the first film directed by [person]
T20 What is the occupation of [person]?
Table 3: Question templates used to create UnseenQA
Appendix B LLMs used in this project
Table 4 summarizes the LLMs used in our experiments.
Models #Params Type Open Source Fine-Tuning Release Date Pre-Training
IT RLHF Knowledge # Token Vocab
[…]
Table 4: Summary of LLMs used in our experiments. ‘IT’ denotes Instruction Tuning, and ‘RLHF’ refers to Reinforcement Learning from Human Feedback. ‘Knowledge’ indicates the knowledge cutoff date. Underlined dates were not explicitly provided by the authors but extrapolated from the datasets used for LLM training. Flan-T5’s base model is T5 version 1.1 pre-trained on the C4 dataset, filtered from web-extracted text in April 2019. Llama 1’s pre-training data includes Wikipedia dumps from June to August 2022.
Appendix C Prompts used in our Experiments
The prompt used to detect time-sensitive questions is shown in Table 5. The QA prompts under three different prompt settings are shown in Table 6. The prompt used to check whether an LLM’s response matches the ground-truth is shown in Table 7. The prompt used to generate distractors for the consistency test is shown in Table 8.
Prompt for detecting time-sensitive questions
INSTRUCTION: Please provide the index of questions whose answers change yearly. Just return the index without explanations.
Here is the list of questions:
1. Who is the most paid player in EPL?
2. What is the capital of Louisiana?
3. Who won the Nobel Peace Prize in 2009?
4. What is the latest model of the iPhone currently available?
Index:
1, 4
Here is the list of questions:
[question placeholder]
Index:
Table 5: The prompt for detecting time-sensitive questions
QA prompt in zero-shot
INSTRUCTION: Please answer knowledge-related questions directly. Note: Please do not give anything other than the answer; Say "unsure" if you do not know.
QUESTION: [question placeholder]
ANSWER:
QA prompt in four-shot
INSTRUCTION: Please answer knowledge-related questions directly. Note: Please do not give anything other than the answer; Say "unsure" if you do not know.
QUESTION: [question example 1 from $R_{seen}$]
ANSWER: [answer 1]
QUESTION: [question example 2 from $R_{seen}$]
ANSWER: [answer 2]
QUESTION: [question example 3 from $R_{seen}$]
ANSWER: [answer 3]
QUESTION: [question example 4 from $R_{seen}$]
ANSWER: [answer 4]
QUESTION: [question placeholder]
ANSWER:
QA prompt in four-shot with tew unsure shot
INSTRUCTION: Please answer knowledge-related questions directly. Note: Please do not give anything other than the answer; Say "unsure" if you do not know.
QUESTION: [question example 1 from $R_{seen}$]
ANSWER: [answer 1]
QUESTION: [question example 2 from $R_{seen}$]
ANSWER: [answer 2]
QUESTION: [question example 3 from $R_{unseen}$]
ANSWER: unsure
QUESTION: [question example 4 from $R_{unseen}$]
ANSWER: unsure
QUESTION: [question placeholder]
ANSWER:
Table 6: The question answering prompt format. The shots are selected from repositories, $R_{seen}$ and $R_{seen}$. The order of shots is random. For the MCQ tests in consistency experiments, we edit the instruction line to INSTRUCTION: Please answer knowledge-related multi-choice questions directly. Note: Please do not give anything other than the appropriate option (A, B, C, D or E); choose the option indicating "unsure" if you do not know.
Prompt for check whether an answer matches the ground truth for the question
INSTRUCTION: You need to check whether the prediction of a question-answering system to a question is correct. You should make the judgment based on a list of ground truth answers provided to you. Your response should be "yes" if the prediction is correct or "no" if the prediction is wrong.
Question: Who authored The Taming of the Shrew (published in 2002)?
Ground truth: ["William Shakespeare", "Roma Gill"]
prediction: W Shakespeare
Correctness: yes
Question: What country is Maharashtra Metro Rail Corporation Limited located in?
Ground truth: ["India"]
prediction: Maharashtra
Correctness: no
Question: Edward Tise (known for Full Metal Jacket (1987)) is in what department?
Ground truth: ["sound department"]
Prediction: 2nd Infantry Division, United States Army
Correctness: no
Question: Which era did Michael Oakeshott belong to?
Ground truth: ["20th-century philosophy"]
prediction: 20th century.
Correctness: yes
Question: [quesetion placeholder]
Ground truth: [ground truth placeholder]
prediction: [LLM’s answer placeholder]
Correctness:
Table 7: The prompt used to check whether an LLM’s answer matches the ground truth for the question
Prompt for generating distractors
INSTRUCTION: For the given question-answer pair, provide 20 different distractors that are similar yet distinct from the given answer. Note: Seperate the 20 distractors with a special token "[SEP]".
Q: Who was the President of the United States in 2010?
A: Barack Obama
Distractors: George W. Bush [SEP] Bill Clinton [SEP] Ronald Reagan [SEP] Donald Trump [SEP] Jimmy Carter [SEP] George H.W. Bush [SEP] Richard Nixon [SEP] Gerald Ford [SEP] Lyndon B. Johnson [SEP] John F. Kennedy [SEP] Dwight D. Eisenhower [SEP] Harry S. Truman [SEP] Franklin D. Roosevelt [SEP] Herbert Hoover [SEP] Calvin Coolidge [SEP] Woodrow Wilson [SEP] William Howard Taft [SEP] Theodore Roosevelt [SEP] William McKinley [SEP] Grover Cleveland
Q: What is the name of the first cloned sheep?
A: the first cloned sheep is dolly.
Distractors: the first cloned sheep is Polly [SEP] the first cloned sheep is Molly [SEP] the first cloned sheep is Holly [SEP] the first cloned sheep is Bella [SEP] the first cloned sheep is Daisy [SEP] the first cloned sheep is Lily [SEP] the first cloned sheep is Rosie [SEP] the first cloned sheep is Millie [SEP] the first cloned sheep is Ellie [SEP] the first cloned sheep is Sally [SEP] the first cloned sheep is Tilly [SEP] the first cloned sheep is Nelly [SEP] the first cloned sheep is Jolly [SEP] the first cloned sheep is Betty [SEP] the first cloned sheep is Annie [SEP] the first cloned sheep is Lucy [SEP] the first cloned sheep is Maggie [SEP] the first cloned sheep is Cindy [SEP] the first cloned sheep is Penny [SEP] the first cloned sheep is Ginny
Q: [QUESTION]
A: [ANSWER]
Distractors:
Table 8: The prompt used to generate distractors for consistency tests.
Appendix D Full Experimental Results
Table 9, Table 10, and Table 11 provides the detailed results of LLMs’ performance on factuality, consitency, and reliablity respectively. Figure 12 shows the rankings of LLMs based on different metrics. Figure 13 compare different LLMs’ factuality, consistency, and reliability performance.
Model Params zero-shot four-shot four-shot-2
Seen Uneen Seen Uneen Seen Uneen
WR ($\downarrow$) CR ($\uparrow$) NCR ($\uparrow$) UR ($\uparrow$) WR ($\downarrow$) CR ($\uparrow$) NCR ($\uparrow$) UR ($\uparrow$) WR ($\downarrow$) CR ($\uparrow$) NCR ($\uparrow$) UR ($\uparrow$)
GPT-3.5 Turbo Unknown 30.40 60.73 30.33 81.70 28.97 61.17 32.20 94.70 27.93 57.30 29.37 99.37
Flan-T5 0.08B 73.03 1.83 -71.20 8.40 85.53 1.63 -83.90 2.77 82.57 1.43 -81.13 34.63
0.25B 82.77 5.47 -77.30 5.50 86.43 5.27 -81.17 2.70 32.67 2.23 -30.43 74.67
0.78B 3.70 2.13 -1.57 100.00 37.80 8.00 -29.80 85.30 9.60 4.17 -5.43 99.90
3B 9.70 7.50 -2.20 98.73 46.00 13.67 -32.33 76.27 23.03 11.23 -11.80 99.73
11B 40.97 20.77 -20.20 67.57 60.63 22.23 -38.40 45.27 42.37 20.20 -22.17 90.00
Llama 1 7B 43.93 35.67 -8.27 23.40 54.47 42.10 -12.37 4.73 27.57 28.27 0.70 54.50
13B 34.33 41.13 6.80 24.63 49.67 47.43 -2.23 4.10 27.37 35.53 8.17 69.37
65B 28.40 46.87 18.47 37.77 39.77 57.73 17.97 19.47 14.63 34.87 20.23 90.10
Llama 2 7B 23.73 31.50 7.77 67.30 54.63 42.03 -12.60 5.33 35.53 30.10 -5.43 66.60
13B 23.67 41.07 17.40 42.83 48.20 49.17 0.97 6.03 27.40 39.07 11.67 71.23
70B 37.17 55.33 18.17 18.17 38.20 59.83 21.63 13.03 19.50 46.03 26.53 95.10
Llama2chat 7B 47.37 36.33 -11.03 60.30 26.87 27.90 1.03 98.60 23.43 23.33 -0.10 99.73
13B 37.87 41.27 3.40 71.30 25.03 39.13 14.10 96.13 32.00 41.27 9.27 94.90
70B 29.53 47.50 17.97 98.10 14.57 34.10 19.53 99.63 15.13 32.60 17.47 100.00
Llama3 8B 50.20 45.13 -5.07 10.73 49.27 48.23 -1.03 6.50 32.77 36.33 3.57 89.03
70B 27.87 55.53 27.67 32.13 33.70 63.60 29.90 25.93 18.87 53.60 34.73 87.63
Llama3Instruct 8B 53.93 42.03 -11.90 79.97 54.00 39.27 -14.73 69.43 54.60 38.73 -15.87 78.73
70B 36.80 59.03 22.23 70.03 38.40 58.10 19.70 68.47 38.90 56.80 17.90 88.60
Mistral 7B 24.00 39.47 15.47 48.13 50.13 47.07 -3.07 13.57 24.70 36.37 11.67 81.73
Mistral-Instruct 7B 44.77 29.90 -14.87 76.50 39.53 28.63 -10.90 93.80 46.63 29.47 -17.17 79.13
Gemma 2B 51.20 24.63 -26.57 28.17 69.17 27.07 -42.10 2.77 39.37 18.67 -20.70 56.07
7B 38.80 39.73 0.93 12.70 56.50 40.53 -15.97 8.67 30.77 31.23 0.47 68.93
Gemma-Instruct 2B 53.80 9.27 -44.53 88.60 13.27 4.30 -8.97 99.93 14.40 3.77 -10.63 99.30
7B 37.13 19.03 -18.10 98.60 16.17 14.20 -1.97 99.97 19.13 13.60 -5.53 99.93
Phi2 3B 65.97 21.43 -44.53 13.83 72.10 21.40 -50.70 14.77 62.07 19.33 -42.73 50.20
Table 9: Factuality performance.
Model Params zero-shot four-shot four-shot-2
$C_{\textit{wrong}}^{s}$($\downarrow$) $C_{\textit{wrong}}^{u}$($\downarrow$) $C_{\textit{wrong}}$($\downarrow$) $C_{\textit{correct}}$ ($\uparrow$) $C_{\textit{wrong}}^{s}$($\downarrow$) $C_{\textit{wrong}}^{u}$($\downarrow$) $C_{\textit{wrong}}$ ($\downarrow$) $C_{\textit{correct}}$ ($\uparrow$) $C_{\textit{wrong}}^{s}$($\downarrow$) $C_{\textit{wrong}}^{u}$($\downarrow$) $C_{\textit{wrong}}$ ($\downarrow$) $C_{\textit{correct}}$ ($\uparrow$)
GPT-3.5 Turbo - 61.79 23.65 42.72 87.10 57.43 19.62 38.53 85.16 48.56 33.68 41.12 79.68
Flan-T5 0.08B 14.49 20.56 17.53 28.64 16.53 25.62 21.07 47.45 18.44 25.87 22.15 47.21
0.25B 35.33 26.31 30.82 62.29 33.36 25.44 29.40 69.40 35.34 22.80 29.07 75.90
0.78B 45.68 - 45.68 85.23 34.16 33.72 33.94 75.12 42.99 35.00 38.99 82.64
3B 45.03 25.26 35.15 84.20 33.07 16.05 24.56 76.85 37.13 35.62 36.38 80.96
11B 41.62 15.38 28.50 80.43 36.40 16.40 26.40 79.84 40.74 15.07 27.90 80.61
Llama 1 7B 25.01 21.70 23.36 37.43 25.37 23.10 24.23 39.65 23.07 20.89 21.98 34.17
13B 35.13 16.41 25.77 59.11 45.60 36.20 40.90 72.49 48.54 25.45 36.99 73.74
65B 58.06 33.84 45.95 83.38 58.35 37.12 47.73 82.38 63.63 16.63 40.13 83.64
Llama 2 7B 26.68 9.37 18.03 50.02 41.51 34.56 38.03 67.07 41.94 25.26 33.60 67.29
13B 62.02 35.69 48.86 83.08 56.12 45.05 50.58 82.05 58.55 31.89 45.22 83.65
70B 63.13 37.52 50.33 84.36 62.07 35.37 48.72 85.06 52.84 6.43 29.63 79.10
Llama2chat 7B 43.66 15.63 29.64 61.23 17.69 12.50 15.09 20.15 19.82 20.62 20.22 17.99
13B 55.79 32.88 44.33 74.62 56.11 28.97 42.54 76.92 52.23 27.39 39.81 77.52
70B 73.61 59.65 66.63 88.71 71.28 30.91 51.10 82.45 67.82 - 67.82 81.88
Llama3 8B 64.17 48.52 56.35 86.50 57.43 35.72 46.58 85.85 37.86 9.51 23.69 78.80
70B 76.82 41.82 59.32 92.86 75.47 41.62 58.55 92.00 64.67 9.43 37.05 86.07
Llama3Instruct 8B 53.08 29.74 41.41 88.86 50.43 13.25 31.84 85.51 37.34 3.80 20.57 79.64
70B 78.14 59.26 68.70 94.24 74.80 43.32 59.06 93.47 67.25 28.17 47.71 93.03
Mistral 7B 56.79 26.70 41.75 84.15 55.94 37.50 46.72 84.87 51.19 13.06 32.13 83.21
Mistral-Instruct 7B 65.84 31.48 48.66 86.09 62.93 30.99 46.96 84.92 61.64 27.63 44.63 84.21
Gemma 2B 30.18 26.26 28.22 37.77 26.93 27.99 27.46 45.90 28.41 22.11 25.26 47.42
7B 62.41 47.24 54.83 86.17 53.39 41.94 47.66 84.22 52.49 22.35 37.42 85.89
Gemma-Instruct 2B 51.65 42.72 47.19 59.64 53.23 15.00 34.11 54.11 50.51 44.52 47.52 49.51
7B 81.92 51.79 66.85 92.43 72.34 30.00 51.17 84.64 80.70 30.00 55.35 90.56
Phi2 3B 30.09 18.27 24.18 54.92 37.21 19.02 28.11 67.34 38.48 15.78 27.13 68.16
Table 10: Consistency performance.
Model Params zero-shot four-shot four-shot-2
Seen Uneen Seen Uneen Seen Uneen
CWR ($\downarrow$) CCR ($\uparrow$) NCCR ($\uparrow$) IUR ($\uparrow$) CWR ($\downarrow$) CCR ($\uparrow$) NCCR ($\uparrow$) IUR ($\uparrow$) CWR ($\downarrow$) CCR ($\uparrow$) NCCR ($\uparrow$) IUR ($\uparrow$)
GPT-3.5 Turbo - 18.78 52.90 34.11 95.67 16.64 52.09 35.45 98.96 13.56 45.66 32.09 99.79
Flan-T5 0.08B 10.58 0.52 -10.06 81.17 14.13 0.77 -13.36 75.09 15.22 0.68 -14.55 83.09
0.25B 29.25 3.41 -25.84 75.13 28.83 3.66 -25.17 75.25 11.54 1.69 -9.85 94.22
0.78B 1.69 1.82 0.13 100.00 12.91 6.01 -6.90 95.04 4.13 3.45 -0.68 99.97
3B 4.37 6.32 1.95 99.68 15.21 10.51 -4.70 96.19 8.55 9.09 0.54 99.90
11B 17.05 16.70 -0.35 95.01 22.07 17.75 -4.32 91.02 17.26 16.28 -0.98 98.49
Llama 1 7B 10.99 13.35 2.36 83.38 13.82 16.69 2.88 77.99 6.36 9.66 3.30 90.50
13B 12.06 24.31 12.25 87.63 22.65 34.38 11.73 65.29 13.28 26.20 12.91 92.20
65B 16.49 39.08 22.59 78.94 23.21 47.56 24.35 70.10 9.31 29.16 19.85 98.35
Llama 2 7B 6.37 15.76 9.39 96.94 22.68 28.19 5.51 67.28 14.90 20.25 5.35 91.56
13B 14.68 34.12 19.44 79.60 27.05 40.34 13.29 57.67 16.04 32.68 16.64 90.83
70B 23.47 46.68 23.21 69.30 23.71 50.89 27.18 69.24 10.30 36.41 26.11 99.69
Llama2chat 7B 20.68 22.24 1.56 93.80 4.75 5.62 0.87 99.83 4.64 4.20 -0.45 99.94
13B 21.13 30.80 9.67 90.56 14.04 30.10 16.05 98.88 16.71 31.99 15.28 98.60
70B 21.74 42.14 20.40 98.87 10.39 28.12 17.73 99.89 10.26 26.69 16.43 100.00
Llama3 8B 32.22 39.04 6.82 56.69 28.30 41.41 13.11 66.60 12.41 28.63 16.22 98.96
70B 21.41 51.57 30.15 71.62 25.43 58.52 33.08 69.17 12.20 46.13 33.93 98.83
Llama3Instruct 8B 28.62 37.35 8.72 94.04 27.23 33.58 6.34 95.95 20.39 30.85 10.46 99.19
70B 28.76 55.63 26.88 82.24 28.72 54.31 25.59 86.34 26.16 52.84 26.68 96.79
Mistral 7B 13.63 33.21 19.58 86.15 28.04 39.95 11.90 67.59 12.65 30.26 17.62 97.61
Mistral-Instruct 7B 29.48 25.74 -3.74 92.60 24.88 24.31 -0.56 98.08 28.74 24.82 -3.92 94.24
Gemma 2B 15.45 9.30 -6.15 81.14 18.62 12.42 -6.20 72.78 11.19 8.85 -2.33 90.29
7B 24.22 34.24 10.02 58.76 30.16 34.14 3.97 61.70 16.15 26.82 10.67 93.06
Gemma-Instruct 2B 27.79 5.53 -22.26 95.13 7.06 2.33 -4.74 99.99 7.27 1.87 -5.41 99.69
7B 30.42 17.59 -12.83 99.27 4.46 12.02 7.55 99.99 15.44 12.32 -3.12 99.98
Phi2 3B 19.85 11.77 -8.08 84.25 26.83 14.41 -12.42 83.79 23.88 13.17 -10.71 92.14
Table 11: Reliability performance.
Figure 12: Ranking of LLMs based on different Metrics.
Figure 13: Visualization of LLMs’ factuality, consistency and reliablity performance.

### v2

@@
-Large Language Models as Reliable Knowledge Bases?
+How Reliable are LLMs as Knowledge Bases? Re-thinking Facutality and Consistency
 Abstract
-The NLP community has recently shown a growing interest in leveraging Large Language Models (LLMs) for knowledge-intensive tasks, viewing LLMs as potential knowledge bases (KBs).
-However, the reliability and extent to which LLMs can function as KBs remain under-explored.
-While previous studies suggest LLMs can encode knowledge within their parameters, the amount of parametric knowledge alone is not sufficient to evaluate their effectiveness as KBs.
-This study defines criteria that a reliable LLM-as-KB should meet, focusing on factuality and consistency, and covering both seen and unseen knowledge.
-[^1] Seen knowledge refers to knowledge learned during training.
-Unseen knowledge is neither present in the model’s training data nor can be inferred from seen knowledge.
-We develop several metrics based on these criteria and use them to evaluate 26 popular LLMs, while providing a comprehensive analysis of the effects of model size, instruction tuning, and in-context learning (ICL).
-Our results paint a worrying picture.
-Even a high-performant model like GPT-3.5-turbo is not factual or consistent, and strategies like ICL and fine-tuning are unsuccessful at making LLMs better KBs.
+Large Language Models (LLMs) are increasingly explored as knowledge bases (KBs), yet current evaluation methods focus too narrowly on knowledge retention, overlooking other crucial criteria for reliable performance.
+In this work, we rethink the requirements for evaluating reliable LLM-as-KB usage and highlight two essential factors: factuality, ensuring accurate responses to seen and unseen knowledge [^1] Seen knowledge refers to knowledge learned during training.
+Unseen knowledge is neither present in the model’s training data nor can be inferred from seen knowledge., and consistency, maintaining stable answers to questions about the same knowledge.
+We introduce UnseenQA, a dataset designed to assess LLM performance on unseen knowledge, and propose new criteria and metrics to quantify factuality and consistency, leading to a final reliability score.
+Our experiments on 26 LLMs reveal several challenges regarding their use as KBs, underscoring the need for more principled and comprehensive evaluation.
+How Reliable are LLMs as Knowledge Bases? Re-thinking
+Facutality and Consistency
+Danna Zheng1, Mirella Lapata1, Jeff Z.
+Pan1,2 1 School of Informatics, University of Edinburgh, UK 2 Huawei Edinburgh Research Centre, CSI, UK dzheng@ed.ac.uk, mlap@inf.ed.ac.uk, http://knowledge-representation.org/j.z.pan/
 1 Introduction
-Large language models (LLMs), pretrained on extensive text corpora, have demonstrated the ability to implicitly encode various types of knowledge within their weights, without requiring human supervision.
-As a result, many recent studies Chuang et al.
-(2023); Yu et al.
-(2023); Dhingra et al.
-(2022); Sung et al.
-(2021); Wang et al.
-(2020) aim to analyze the relationship between LLMs and KBs, and even explore whether LLMs can replace KBs Sun et al.
-(2023); Mruthyunjaya et al.
-(2023); Heinzerling and Inui (2021).
-However, whether current LLMs can serve as reliable KBs and how to evaluate their performance in this role remains largely unexplored.
-Existing studies Sun et al.
+Large Language Models (LLMs), pretrained on vast text corpora, have demonstrated significant capabilities in encoding knowledge without explicit supervision.
+The continuous release of new LLMs, evaluated on benchmarks like TriviaQA Joshi et al.
+(2017) and Natural Questions Kwiatkowski et al.
+(2019), highlights their improving ability to answer fact-based queries.
+This progress has fueled interest in employing LLMs as knowledge bases (KBs) for various applications and developing techniques to edit model knowledge Wang et al.
+(2024c, b, a) or mitigate hallucinations Zhang et al.
+(2024b, a); Yu et al.
+(2024).
+However, a critical question remains underexplored: What criteria should an LLM meet to function reliably as a KB?
+Current research often assumes that knowledge retention alone is sufficient Sun et al.
 (2023); Wang et al.
 (2021); Roberts et al.
-(2020) often implicitly assume that the LLM’s ability to retain knowledge is sufficient for it to function as a KB.
-Typically, these studies employ two methods: (1) converting knowledge graphs into natural language questions using templates and evaluating LLM ability to answer these questions, by measuring the amount of knowledge therein Petroni et al.
+(2020).
+Existing evaluations generally follow two approaches: (1) converting knowledge graphs into natural language questions and assessing how many questions the LLM answers correctly Petroni et al.
 (2019); Sun et al.
-(2023); and (2) pre-training LLMs on passages/triples containing knowledge and then assessing their ability to answer related questions, thereby quantifying their knowledge retention Wang et al.
+(2023); and (2) pretraining LLMs on knowledge-rich text and measuring their accuracy on related questions Wang et al.
 (2021); He et al.
 (2024).
-While these methods demonstrate that LLMs can recall knowledge, memorizing facts is not the sole criterion for being a reliable KB AlKhamissi et al.
+Figure 1: (a) An example illustrating three answer types: correct, wrong, and uninformative.
+Focusing only on the correct rate incorrectly suggests that Model B is better, even though Model A is more reliable with a similar correct rate and a much lower wrong rate.
+(b) Illustration of LLM inconsistency with davinci-002 (temperature is set to 0).
+Questions in the top focus on seen knowledge, with probability distribution mass concentrated on one prediction.
+Questions in the bottom focus on unseen knowledge, where the distribution is more even.
+Drawing from such a distribution inevitably leads to inconsistencies.
+(c) Example computation for consistency score $Cons(q,r)$.
+The LLM’s original answer is shown in green, while distractors are red.
+(d) An example illustrating how to evaluate LLM-as-KB.
+These methods demonstrate that LLMs can recall information, but knowledge volume alone does not guarantee reliable performance as a KB.
+Beyond retention, it is essential to examine how LLMs handle factual queries—specifically, whether they respond accurately to seen knowledge and avoid making claims about unseen knowledge (factuality), and whether they provide consistent answers to questions about the same facts (consistency).
+Factuality refers to the quality of being factual or based on fact.
+KBs hosted on servers or cloud platforms, offer precise answers or null responses when data is unavailable.
+In contrast, LLMs rely on probabilistic next-token prediction, which can lead to plausible but incorrect answers.
+As a result, LLM responses are typically correct, uninformative, or wrong.
+Existing methods for evaluating factuality often focus on the rate of correct answers in factual QA datasets Chen et al.
+(2023); Wang et al.
+(2024d).
+However, as many studies Lin et al.
+(2022); Sun et al.
+(2023) fail to specify whether the dataset’s knowledge was included in the LLM’s pre-training data, it is not possible to establish whether the model is genuinely factual.
+Secondly, it is misleading to equate a higher correct rate with greater factuality.
+As illustrated in Figure 1(a), a model with a higher correct rate might still be less factual if it produces a higher rate of wrong answers compared to a model with fewer errors overall.
+Consistency refers to the quality of always behaving in the same way or having the same opinions.
+Traditional KBs achieve consistency through algorithms Andersen and Pretolani (2001) that detect and resolve conflicts.
+In contrast, LLMs frequently exhibit inconsistent behavior Elazar et al.
+(2021); Wang et al.
 (2022).
-What criteria then should a LLM meet to function as a reliable KB?
-Discussion on this topic has been limited, and there is no agreement on the definition of these criteria.
-AlKhamissi et al.
-(2022) argue that LLMs ought to excel at five aspects (i.e., access, edit, consistency, reasoning, explainability, and interpretability) if they are to be considered a KB.
-However, they do not outline specific metrics to evaluate the extent to which LLMs act as KBs.
-Besides, we argue that evaluating LLMs against the characteristics of KBs may not be entirely appropriate due to their different data storage structures.
-Instead, we should consider the specific properties of LLMs when assessing their suitability as KBs.
-Our research addresses these gaps and aims to establish a more nuanced understanding of LLMs-as-KBs.
-We evaluate and compare the reliability of different LLMs functioning as KBs in answering factoid questions.
-Specifically, our work makes the following contributions:
-  1. 1.
-    We define how to assess the reliability of LLMs-as-KBs, and propose metrics that consider the differences between KBs and LLMs and align with the distinct nature of LLMs.
-We consider two aspects, factuality (i.e., the ability to provide factual responses) and consistency (i.e., the ability to provide consistent responses for questions involving the same knowledge), and cover the evaluation on both seen and unseen knowledge.
-  2. 2.
-    To evaluate consistency, we propose a novel method which computes the probability that a LLM can consistently provide response $r$ for question $q$.
-  3. 3.
-    To evaluate LLMs on unseen knowledge, we create a new QA dataset UnseenQA, with a knowledge cutoff date before April 13, 2024.
-  4. 4.
-    We evaluate 26 popular LLMs in their ability to function as reliable KBs, and discuss the influence of model size, instruction-tuning, and ICL on their performance.
-Figure 1: An example illustrating three answer types: correct, wrong, and uninformative.
-Focusing only on the correct rate incorrectly suggests that Model B is better, even though Model A is more reliable with a similar correct rate and a much lower wrong rate.
+Current research Elazar et al.
+(2021); Jang et al.
+(2022); Hagström et al.
+(2023) evaluates LLM consistency using benchmarks involving paraphrasing, negation, or multilingual variations, favoring models that maintain consistent responses across diverse samples.
+However, we argue that expecting LLMs to always be consistent in fact-based responses is overly rigid.
+Unlike KBs, which store information in fixed locations, LLMs operate probabilistically.
+When the context has been learned during training, the probability distribution for predictions is concentrated; otherwise, it remains more uniform.
+Sampling from a uniform distribution naturally leads to inconsistencies.
+As shown in Figure 1(b), even with greedy decoding, slight distribution biases can cause variations in the top-selected words.
+Given these issues, this paper seeks to define the criteria for a reliable LLM-as-KB when handling factual queries.
+In evaluating factuality, we consider both seen knowledge (contained within training data) and unseen knowledge, and consider the negative effects of wrong answers.
+To assess performance on unseen knowledge, we introduce UnseenQA, a new dataset containing knowledge unavailable to LLMs trained before April 13, 2024.
+For consistency, we classify the correctness of responses and propose a novel method to compute the probability that an LLM can consistently provide the same response $r$ to a question $q$.
+We evaluate 26 popular LLMs and find that: 1) GPT-3.5-Turbo achieves balanced performance in both factuality and consistency.
+LLMs like LLaMA3-70B may achieve a higher correct rate on seen knowledge but exhibit higher wrong rates and consistency on wrong answers.
+2) There is a correlation between factuality and consistency: more factual LLMs tend to be consistent in their responses, whether correct or wrong.
+3) Larger LLMs perform worse on unseen knowledge and are more consistent even when providing wrong answers.
+4) Fine-tuning techniques can improve performance on unseen knowledge.
+However, this often comes at the expense of performance on seen knowledge.
+5) In-context learning (ICL) does not improve performance on seen knowledge, as it as it generally increases or decreases both correct and wrong rates simultaneously.
+6) Base LLMs tend to overestimate their knowledge on numerical and temporal questions.
+We hope our work will draw the community’s attention to the multifaceted challenges of using LLMs-as-KBs, ultimately inspiring further research and innovation toward more reliable, robust, and principled methodologies.
 2 What is a Reliable LLM-as-KB?
-In simple terms, a LLM functions as a reliable KB if it consistently provides factual responses.
-Evaluating the reliability of LLMs as KBs primarily involves assessing two critical dimensions, namely factuality and consistency.
+In simple terms, an LLM is a reliable KB if it consistently provides factual responses.
+As illustrating in Figure 1 (d), evaluating the reliability of LLMs as KBs primarily involves assessing two critical dimensions, namely factuality and consistency.
 2.1 Factuality
-Factuality refers to the quality of being factual or based on fact.
-Run-of-the-mill KBs, stored on physical servers or cloud platforms, deliver information directly in response to queries.
-If the requested data is unavailable, these systems typically return a null response.
-In contrast, LLMs are probabilistic models that excel at next word prediction based on the given context, rather than storing explicit information in defined locations.
-This architecture allows LLMs to generate responses that seem plausible, regardless of whether the content was included in their training data.
-Consequently, LLMs typically produce three types of responses: correct, uninformative, and wrong.
-Ongoing efforts Chen et al.
-(2023); Wang et al.
-(2024) to evaluate the factuality of LLMs often hinge on the models’ correct rates in responding to factual QA datasets.
-However, this approach has notable limitations.
-Firstly, many studies Lin et al.
-(2022); Sun et al.
-(2023) do not specify whether the dataset’s scope of knowledge was included in the LLM’s training data.
-This omission can lead to unfair comparisons between models, especially if the knowledge being tested is within the training scope of one model but not another.
-Secondly, the assumption that a higher correct rate indicates greater factuality is problematic.
-For example, consider models A and B shown in Figure 1, and assume they are being evaluated on a test dataset covering knowledge they both have seen during training.
-In this case, focusing solely on the correct rate might erroneously suggest that Model B is more factual, despite Model A being more reliable due to its similar correct rate and significantly lower wrong rate.
-Given these issues, we propose the following criteria for the factuality of LLMs-as-KBs:
-CRITERION 1.1
-For seen knowledge, a factual LLM should demonstrate a high correct rate and a low wrong rate.
-CRITERION 1.2
-For unseen knowledge, a factual LLM should demonstrate a high uninformative rate.
+We propose the following criteria for determining the factuality of LLMs-as-KBs:
+Criterion 1.1: For seen knowledge, a factual LLM should demonstrate a high correct rate and a low wrong rate.
+Criterion 1.2: For unseen knowledge, a factual LLM should demonstrate a high uninformative rate.
 We next proceed to define evaluation metrics that operationalize these criteria.
-Let $\boldsymbol{M}$ denote a LLM.
+Let $\boldsymbol{M}$ denote an LLM.
 Let $D_{\text{seen}}$ denote a QA dataset containing $N$ open-ended factoid questions pertaining to knowledge the LLM ought to have seen during training.
@@
 A higher UR indicates that the model is more likely to refrain from giving wrong responses when faced with unseen knowledge.
-Figure 2: Illustration of LLM inconsistency with davinci-002 (temperature is set to 0).
-Questions in the top focus on seen knowledge, with probability distribution mass concentrated on one prediction.
-Questions in the bottom focus on unseen knowledge, where the distribution is more even.
-Drawing from such a distribution inevitably leads to inconsistencies.
 2.2 Consistency
-Consistency refers to the quality of always behaving in the same way or having the same opinions.
-KBs are designed with this principle in mind.
-In fact, there are efficient algorithms Andersen and Pretolani (2001) which detect and resolve conflicts within KBs, thus ensuring consistent outputs for queries on the same facts.
-It is well-known that LLMs often exhibit inconsistencies in their responses Elazar et al.
-(2021); Wang et al.
-(2022).
-Current research Elazar et al.
-(2021); Jang et al.
-(2022); Hagström et al.
-(2023) evaluates LLM consistency through their performance on benchmarks involving paraphrasing, negation, or multilingual variations.
-A model is considered superior if it responds consistently across a broader range of data samples.
-In this work, we argue that it may be too strict to expect LLMs to be always consistent when responding to fact-based questions.
-Unlike explicit KBs which store information at a fixed location, LLMs operate probabilistically.
-In theory, if the context has been learned during training, the probability distribution for the prediction will be concentrated; otherwise, it will be more uniform.
-Drawing from a uniform probability distribution inevitably leads to inconsistencies.
-As shown in Figure 2, even with greedy decoding, slight biases in the distribution can cause fluctuations in the selection of the top probable words.
-Given their probabilistic nature, we do not expect LLMs to always behave consistently.
-We acknowledge that inconsistencies can cause confusion in practical applications and propose to monitor model behavior through post-processing which we argue is more realistic than expecting a probabilistic model to be perpetually consistent.
-We thus propose the following consistency criteria:
-CRITERION 2.1
-The model is expected to be consistent when it produces correct responses.
-CRITERION 2.2
-The model is expected to be inconsistent when it produces wrong responses.
+We propose the following consistency criteria:
+Criterion 2.1: The model is expected to be consistent in correct responses. Criterion 2.2: The model is expected to be inconsistent in wrong responses.
 We next define evaluation metrics corresponding to the criteria above.
@@
 (2024), we measure consistency based on multiple-choice questions (MCQs).
-As shown in Figure 3, we employ gpt-3.5-turbo-insruct to generate a set of distractor options similar to response $r$, and then create a group of MCQs.
+As shown in Figure 1 (c), we employ gpt-3.5-turbo-insruct to generate a set of distractor options similar to response $r$, and then create a group of MCQs.
 The consistency score for data point $(q,r)$ is calculated as:
-$Cons(q,r)=\frac{\sum_{i=1}^{X_{\text{MCQs}}}[R_{i}=r]}{X_{\text{MCQs}}}$ (4)
-where $X_{\text{MCQs}}$ is the total number of MCQs, $R_{i}$ is model $M$’s response for the $i$-th MCQ, and the expression $[R_{i}=r]$ yields 1 when the model’s response $R_{i}$ matches its original response $r$, and 0 otherwise.
-The consistency score $Cons(q,r)$ ranges from 0 to 1.
+$\text{Cons}(q,r)=\frac{\sum_{i=1}^{X_{\text{MCQs}}}[R_{i}=r]}{X_{\text{MCQs}}}$ (4)
+where $X_{\text{MCQs}}$ is the total number of MCQs, $R_{i}$ is model $M$’s response for the $i$-th MCQ, and $[R_{i}=r]$ yields 1 when the model’s response $R_{i}$ matches its original response $r$, and 0 otherwise.
+The consistency score $\text{Cons}(q,r)$ ranges from 0 to 1.
 METRIC 2.1: $C_{\textit{correct}}$
-measures the consistency of the model when it provides correct responses. It is defined as:
+measures a model’s consistency in its correct responses and is defined as:
 $C_{\textit{correct}}=\frac{\sum_{j=1}^{N_{correct}}Cons(q_{j}^{(c)},r_{j}^{(c)})}{N_{correct}}$ (5)
@@
 METRIC 2.2: $C_{\textit{wrong}}$
-measures the consistency of an LLM when it provides wrong responses. It is defined as:
+measures the consistency of an LLM when it provides wrong responses and is defined as:
 $C_{\textit{wrong}}=\frac{C_{\textit{wrong}}^{s}+C_{\textit{wrong}}^{u}}{2}$ (6)
-where $C_{\textit{wrong}}^{s}$ refers to the consistency of a LLM when it provides wrong responses to questions about seen knowledge, and $C_{\textit{wrong}}^{u}$ refers to the consistency of a LLM when it provides wrong responses to questions about unseen knowledge:
+where $C_{\textit{wrong}}^{s}$/$C_{\textit{wrong}}^{u}$ refer to the consistency of an LLM when it provides wrong responses to questions about seen/unseen knowledge:
 $C_{\textit{wrong}}^{s}=\frac{\sum_{j=1}^{N_{wrong}}Cons(q_{j}^{(w)},r_{j}^{(w)})}{N_{wrong}}$ (7)
 $C_{\textit{wrong}}^{u}=\frac{\sum_{j=1}^{L_{wrong}}Cons(q_{j}^{(w)},r_{j}^{(w)})}{L_{wrong}}$ (8)
-where $r^{(w)}$ refers to responses labeled as wrong, and $q^{(w)}$ is the corresponding question.
-$N_{wrong}$ and $L_{wrong}$ are counts of wrong responses to $D_{\text{seen}}$ and $D_{\text{unseen}}$, respectively.
-$C_{\textit{wrong}}$ ranges from 0 to 1.
-Based on CRITERION 2.2, a lower $C_{\textit{wrong}}$ is desirable.
-Figure 3: Example computation for consistency score $Cons(q,r)$. The LLM’s original answer is shown in green, while distractors are red.
+where $r^{(w)}$ and $q^{(w)}$ denote wrong responses and their corresponding questions.
+$N_{wrong}$ and $L_{wrong}$ are the counts of wrong answers on $D_{\text{seen}}$ and $D_{\text{unseen}}$, respectively.
+$C_{\textit{wrong}}$ ranges from 0 to 1, and per Criterion 2.2, lower $C_{\textit{wrong}}$ is better.
 2.3 Reliability (Factuality and Consistency)
-A reliable LLM-as-KB should then be assessed against factuality and consistency. Based on the criteria defined above, a LLM-as-KB is reliable if it meets the following:
-CRITERION 3.1
-For seen knowledge, a LLM should have a high rate of consistently correct responses and a low rate of consistently wrong responses.
-CRITERION 3.2
-For unseen knowledge, a LLM should have a high rate of uninformative or inconsistent responses.
-We quantify these criteria with two metrics.
+Based on the criteria defined above, an LLM is reliable as a KB if it meets the following criteria when evaluated against factuality and consistency:
+Criterion 3.1: For seen knowledge, an LLM should have a high rate of consistently correct responses and a low rate of consistently wrong responses.
+Criterion 3.2: For unseen knowledge, a LLM should have a high rate of uninformative or inconsistent responses.
+We next quantify these criteria are with the following two metrics.
 METRIC 3.1: Net Consistently Correct Rate (NCCR)
@@
 NCCR ranges from $-1$ to $1$.
-A higher NCCR indicates a LLM is more reliable on seen knowledge.
+NCCR values closer to 1 indicate an LLM is more reliable on seen knowledge.
 A negative NCCR suggests the model provides consistently wrong responses, while a positive NCCR suggests a preference for consistently correct responses.
 METRIC 3.2: Inconsistent/Uninformative Rate (IUR)
-assesses whether the model is likely to provide uninformative or inconsistent wrong responses for questions about unseen knowledge. It is defined as:
+assesses whether an LLM is likely to provide uninformative or inconsistent wrong responses for questions about unseen knowledge. It is defined as:
 $\text{IUR}=1-(1-\text{UR})C_{wrong}^{u}$ (10)
 IUR ranges from 0 to 1. A higher IUR value indicates the LLM functions as a more reliable KB on unseen knowledge.
-LLM Zero-Shot Four-Shot Four-Two
-NCCR IUR AVG NCCR IUR AVG NCCR IUR AVG
-gpt-3.5-turbo 34.1 95.7 62.7 35.5 99.0 66.7 32.1 99.8 65.8
-llama2chat-70b 20.4 98.9 59.1 17.7 99.9 58.8 16.4 100.0 58.2
-llama3-70b 30.2 71.6 36.7 33.1 69.2 35.7 33.9 98.8 65.8
-Table 1: Most reliable LLMs across three prompt settings.
-Four-Two refers to the four-shot with two unsure shots setting.
-All numbers shown are percentages.
-AVG represents average NCCR and (normalized) IUR ($2*\text{IUR}-100$) scores (both scores use the same scale).
-We treat seen and unseen knowledge equally, however, in practice, NCCR and IUR can be assigned different weights.
 3 Experimental Setup
-3.1 Datasets
-To evaluate the performance of LLMs on seen knowledge, we collated SeenQA, a composite dataset containing 3,000 questions sourced from the test sets (or development sets, where test sets were unavailable) of Natural Questions Kwiatkowski et al.
-(2019), TriviaQA Joshi et al.
-(2017), and PopQA Mallen et al.
-(2023).
-For assessing LLM performance on unseen knowledge, we introduce UnseenQA, a new dataset designed to ensure that models with a knowledge cutoff date before April 13, 2024 do not have access to answers.
-UnseenQA also includes 3,000 questions, which are derived from 20 templates covering various answer types such as number, people, time, location, and others.
-Our methodology for creating SeenQA and UnseenQA is detailed in Appendix A.
 3.1 LLM Selection
-We evaluate a wide range of popular LLMs (26 in total) in their ability to function as KBs and investigate the impact of various factors, such as the number of parameters and fine-tuning.
-Specifically, we consider the following models: gpt-3.5-turbo, flan-t5 (0.08B, 0.25B, 0.78B, 3B, 11B), llama1 (7B, 13B, 65B), llama2 (7B, 13B, 70B), llama2-chat (7B, 13B, 70B), llama3 (8B, 70B), llama3instruct (8B, 70B), mistral (7B), mistral-instruct (7B), gemma (2B, 7B), gemma-instruct (2B, 7B), and Phi2 (3B).
-Detailed descriptions of these models are provided in Table 4 in Appendix B.
-We classify LLMs into three categories based on their parameter sizes: small (0.08B–3B), medium (7B–13B), and large (65B–70B).
-We use the term ‘fine-tuned LLMs’ to refer to the LLMs that have been fine-tuned through instruction-tuning or reinforcement learning from human feedback.
-We use the term ‘base LLMs’ to refer to LLMs without fine-tuning.
-3.3 Evaluation on a Single Response
-Uninformative
-We identify three types of uninformative responses from LLMs: ‘repetition’, ‘none’, and ‘unsure’.
-‘Repetition’ refers to responses that repeatedly echo a specific string.
-We detect this using regular expressions and word frequency analysis.
-‘None’ denotes responses lacking relevant information, such as an empty string and repetition of the question.
-‘Unsure’ indicates responses where the model explicitly states it is unable to answer or lacks the required knowledge.
-We label a response as ‘unsure’ if it includes expressions such as ‘I am not sure’, ‘I cannot provide’, and ‘I am just an AI’.
-Correct
-We determine whether a response is correct based on exact match.
-A response is considered correct if the exact match score is 1.
-In cases where exact match is 0, we compare the model’s prediction against the ground truth using gpt-4o.
-Previous work Sun et al.
-(2023) has shown a 98% agreement rate between ChatGPT and human judgments in comparing model responses with ground truth.
-We follow their prompt as outlined in Table 7 in Appendix C.
-Consistency Score
-To compute $Cons(q,r)$, we set $X_{\text{MCQs}}$ (total number of MCQs) to 20, and each MCQ includes question $q$ and 5 options (the original response $r$, 3 random distractor options, and an ‘unsure’ option).
-3.4 Prompts and Hyper-parameters
-To provide a comprehensive evaluation, we experimented with three types of prompt settings: zero-shot, four-shot, and four-shot with two unsure shots.
-To avoid any bias introduced by fixed examples, we employed a dynamic few-shot method following the work of Nori et al.
-(2023).
-We collected two repositories, $R_{seen}$ and $R_{unseen}$.
-$R_{seen}$ includes 280 question-answer pairs about seen knowledge (200 from the unused data of PopQA and training data of Natural Questions and TriviaQA; 80 are generated using the templates in Table 3).
-$R_{unseen}$ consists of 40 question-answer pairs about unseen knowledge, all generated using the templates in Table 3.
-We used text-embedding-3-small to embed the questions in the repositories and test questions as vector representations.
-For each test question under the four-shot setting, we retrieved its nearest four questions from $R_{seen}$.
-Under the four-shot with two unsure shots setting, we retrieved the nearest two questions from $R_{seen}$ and two from $R_{unseen}$.
-All LLMs were evaluated using greedy decoding (temperature 0 for gpt-3.5-turbo) with a maximum of 100 new tokens. Our prompts are provided in Appendix C.
-Figure 4: Comparing llama3instruct-8b against llama2chat-13b, and phi2-3b against gemma-2b using CR, WR, and NCR metrics (zero-shot setting).
-Even though these models are comparable according to CR, they are quite different according to WR and NCR metrics.
-We observe similar trends in the four-shot and with two unsure shots settings.
-4 Results
-We present detailed results for all LLMs are in Table 9 (factuality), Table 10 (consistency), and Table 11 (reliability) in Appendix D.
-LLM rankings based on different metrics are shown in Figure 12 and Figure 13, also in Appendix D.
-gpt-3.5-turbo is overall the most reliable LLM.
-Table 1 presents results for the two best performing LLMs under different prompt settings.
-As can be seen, gpt-3.5-turbo is most reliable across the board.
-Although it is not consistently wrong when asked about facts it does not know (its IUR score exceeds 95%), it is not consistently correct when asked about facts it has seen before (NCCR is only 32%).
-Figure 5: LLM ranking based on NCCR and IUR metrics in zero-shot setting. See full results in Figure 12 (Appendix D) for other shot settings.
-flan-t5-0.78B is the most reliable LLM for unseen knowledge and most unreliable with seen knowledge.
-Figure 5 shows the LLMs ranked by NCCR and IUR in a zero-shot setting.
-As can be seen, while flan-t5-0.78B ranks low for NCCR, it maintains the top position for IUR.
-A similar trend is observed with the flan-t5-3B and gemma-instruct (2B, 7B) models.
-Conversely, models in the llama3 family show the opposite trend: they rank high on seen knowledge but low on unseen knowledge.
-For instance, llama3-70b, despite ranking second in terms of NCCR, falls to the fourth lowest position for IUR.
-Net Correct Rate (NCR) reveals factuality gaps in LLMs.
-As illustrated in Figure 4, CR, the standard metric for assessing factuality, fails to fully capture nuanced differences.
-Despite similar CR values of approximately 40%, llama3instruct-8B and llama2chat-13B behave differently when it comes to wrong responses (on seen data); the former model has a WR of 15% higher than the latter, and as a result its NCR is substantially lower.
-In the case of phi2-3B, Gunasekar et al.
-(2023) claim that with "textbook quality" data, smaller LLMs can achieve satisfactory performance with less training data.
-Focusing solely on CR, this claim is true as phi2-3B’s CR is comparable to that of gemma-2B, which was trained with twice the token count.
-However, phi2-3B exhibits significantly higher WR (and lower NCR) compared to gemma-2B which underscores the challenge of maintaining low error rates even with superior data quality.
-A less factual LLM can be more reliable.
-Figure 6 illustrates this finding.
-The llama3instruct-8b model has a slightly worse NCR compared to llama2chat-7B, indicating that the former model is marginally less factual on seen knowledge.
-However, llama3instruct-8b exhibits a higher NCCR due to its better consistency score on correct responses.
-This suggests that llama3instruct-8b is more reliable than llama2chat-7B on seen knowledge.
-Additionally, both LLMs show negative NCR and positive NCCR in the zero-shot setting.
-This indicates that although they produce more wrong than correct responses, they generate fewer consistently wrong responses compared to consistently correct ones.
-Figure 6: Comparison between llama3instruct-8b and llama2chat-7b (zero-shot setting).
-The former model is slightly less factual according to NCR, but more consistent in its responses (see NCCR, $C_{correct}$, and $C^{s}_{wrong}$).
-We observe similar trends in the four-shot and with two unsure shots settings.
-Comparisons Zero-Shot Four-Shot Four-Two
-NCR vs. UR 0.27 0.34 0.62
-NCCR vs. IUR --0.17 --0.12 0.41
-$C_{\textit{correct}}$ vs. $C_{\textit{wrong}}$  0.81  0.78 0.51
-Table 2: Pairwise correlation of LLM performance on different metrics under different prompt settings.
-The correlation values are computed across all LLMs (26 data points).
-We report Pearson’s $\rho$, diacritic  denotes statistical significance ($p<0.05$).
-Four-Two refers to the four-shot setting with two unsure shots.
-Performance on seen knowledge is not predictive of performance on unseen knowledge.
-In Table 2, we examine whether metrics applied to seen knowledge can be used to extrapolate model performance on unseen knowledge by reporting correlation values (Pearson’s $\rho$) between NCR and UR, and NCCR and IUR.
-As can be seen, in both zero- and four-shot settings, correlations are not statistically significant.
-Nevertheless, correlations are significant in all metric comparisons in the four-shot with two unsure shots setting.
-While in general model performance with seen knowledge does not transfer to unseen knowledge, specific prompt manipulations can enhance the correlation between metrics (see last column in Table 2).
-LLMs are consistently right and wrong!
-Table 2 reports a positive, significant correlation between $C_{correct}$ and $C_{wrong}$.
-This result implies that LLMs demonstrating high consistency in correct responses also tend to exhibit high consistency in wrong responses.
-This finding contradicts our expectation of high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$ and highlights a notable flaw in current models.
-Future work should address this issue, e.g., by instructing LLMs to achieve the desired consistency behavior.
-Figure 7: The impact of model size on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR.
-Different metrics are color-coded.
-LLMs are shown in three sizes, small, medium, and large and are grouped into ‘base’ and fine-tuned ones.
-5 Analysis
-In this section, we explore the effects of model size, fine-tuning, and ICL on LLM performance based on different metrics.
-We also analyze how LLMs handle unseen knowledge by looking at the distribution of uninformative responses and the impact of different question types.
-Figure 8: The impact of fine-tuning on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR.
-Different metrics color-coded.
-This analysis only considers the performance of Llama2, Llama3, Mistral, and Gemma as these families include both base LLMs and fine-tuned versions.
-Models are shown in three sizes, small, medium, and large.
-Figure 9: The impact of ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR.
-Different metrics are color-coded.
-We compare zero-shot and four-shot settings; and zero-shot against four-shot with two unsure shots.
-LLMs (‘base’ and fine-tuned ones) are in three sizes, small, medium, and large.
-Figure 10: The impact of question type on LLM performance as measured by Uninformative Rate (UR) on unseen knowledge.
-Questions are grouped into types based on answer they require.
-Higher values have darker shades.
-5.1 The Impact of Model Size
-Larger LLMs perfom better on seen knowledge but worse on unseen knowledge.
-Figure 7 demonstrates that as model size increases, both NCR (blue line) and NCCR (purple line) improve.
-This trend indicates that larger LLMs perform better on questions about seen knowledge.
-However, for base LLMs in both zero-shot and four-shot settings, IUR (brown line) decreases as model size increases.
-This trend suggests that as LLMs become larger, they become more consistent at delivering wrong responses for questions regarding unseen knowledge.
-Larger LLMs are more consistent, even with wrong responses.
-In Figure 7, we observe that as model size increases, both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) increase significantly.
-While higher consistency scores for correct responses are expected and useful, the higher consistency scores for wrong responses pose a potential risk.
-Larger models may consistently produce convincing but wrong information, which could lead to misinformation if not carefully managed.
-5.2 The Impact of Fine-tuning
-Fine-tuning improves performance on unseen knowledge but negatively affects performance on seen knowledge.
-As illustrated in Figure 8, UR (orange line) and IUR (brown line) show significant improvement in LLMs after fine-tuning, indicating an enhanced ability to distinguish and respond to unseen knowledge appropriately.
-However, the decreasing NCCR (purple line) suggests they become worse at handling seen knowledge after fine-tuning.
-Fine-tuning does not make LLMs more consistent.
-As depicted in Figure 8, there is no noticeable increase in $C_{\textit{correct}}$ (green line) after fine-tuning.
-In fact, fine-tuning even has a negative impact on $C_{\textit{correct}}$ for medium-sized LLMs.
-These results indicate that current fine-tuning techniques fail to enhance the consistency of correct responses.
-Furthermore, $C_{\textit{wrong}}$ (red line) does not decrease after fine-tuning either, which suggests that fine-tuning also fails to reduce the model’s presistence on wrong responses.
-5.3 The Impact of ICL
-Unsure shots improve LLM performance on unseen knowledge.
-As shown in Figure 9, incporating two unsure shots in the four-shot setting substantially increases UR (orange line) and IUR (brown line) across model sizes.
-In contrast, performance on unseen knowledge deteriorates when using four-shots only for all sizes of base LLMs and small fine-tuned LLMs, compared to their zero-shot counterparts.
-ICL does not improve LLM perfomance on seen knowledge.
-Figure 9 demonstrates that ICL does not improve LLM performance according to NCR (blue line) or NCCR (purple line) metrics.
-For small/medium base LLMs and small fine-tuned LLMs in the four-shot setting, ICL even decreases NCR performance.
-This finding contrasts with previous results about the ability of LLMs to learn from a few examples Brown et al.
-(2020); Chada and Natarajan (2021); Touvron et al.
-(2023); Bai et al.
-(2023).
-Earlier work has mostly focused on the correct rate of LLMs in few-shot settings without unsure shots.
-Based on the results presented in Table 9 in Appendix D, we observe that for most LLMs, the four-shot setting indeed substantially improves the correct rate but also increases the wrong rate.
-Incorporating unsure shots helps decrease the wrong rate but also decreases the correct rate.
-ICL makes LLMs less consistent, even on correct responses!
-Figure 9 shows that ICL leads to reductions in both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) for medium/large fine-tuned LLMs and large base LLMs.
-While a lower $C_{\textit{wrong}}$ is desirable, the decrease in $C_{\textit{correct}}$ is concerning.
-This indicates that ICL negatively affects the reliability of LLMs at providing correct responses.
-Figure 11: Distribution of uninformative responses given by LLMs to questions about unseen knowledge.
-We report results for the llama3-8b, gemma-7b, and their fine-tuned models (fourth column) but observe similar trends on other models (omitted for the sake of brevity).
-5.4 Model Behavior on Unseen Knowledge
-Base LLMs overestimate their knowledge on numerical and temporal questions.
-Figure 10 reports UR performance across models, broken down per question type (e.g., number, person, location).
-As can be seen, most base LLMs exhibit significantly lower UR performance on questions requiring numerical or temporal responses.
-This indicates that base LLMs often provide misleading information for such questions, even when they lack the relevant knowledge.
-Fine-tuned LLMs can explicitly admit when they don’t know.
-As illustrated in Figure 11 (fourth column), fine-tuned LLMs are able to explicitly acknowledge their lack of knowledge by answering ‘unsure’ (see definition in Section 3.3) to questions about unseen knowledge.
-In contrast, base LLMs often produce responses classified as ‘none’ or ‘repetition’ in the absence of unsure shots (contrast columns one and two with column three in Figure 11).
-6 Related Work
-Petroni et al.
-(2019) were the first to introduce the concept of using pre-trained language models as knowledge bases.
-They introduced LAMA, a benchmark with questions structured as "fill-in-the-blank" cloze statements, and found that BERT Devlin et al.
-(2019) retains relational knowledge competitive with traditional NLP methods that have some access to oracle knowledge.
-Roberts et al.
-(2020) were the first to measure the extent to which language models trained on unstructured text can implicitly store and retrieve knowledge using natural language queries.
-Specifically, they fine-tuned T5 Raffel et al.
-(2019) to answer questions without access to any external context or knowledge and showed that this approach performs competitively with open-domain systems that explicitly retrieve answers from an external knowledge source when answering questions.
-On a similar vein, Wang et al.
-(2021) fine-tuned BART Lewis et al.
-(2020) with related passages to instill factual knowledge and used the accuracy of masked span recovery to measure the extent to which this knowledge was memorized by the model.
-They found it was challenging for pre-trained language models like BART to remember facts seen in training, and answer questions, even in cases where the relevant knowledge was retained.
-He et al.
-(2024) explicitly trained T5 Raffel et al.
-(2019) and LLAMA2 Touvron et al.
-(2023) to memorize world knowledge from Wikidata on a large scale and then used exact match and F1 scores to evaluate knowledge retention.
-Their results showed LLMs hold promise as large-scale KBs capable of retrieving facts and responding with flexibility, but are less proficient at inferring new knowledge through reasoning.
-Sun et al.
-(2023) proposed a benchmark consisting of 18,000 question-answer pairs representing facts with varying popularity (i.e., high, medium, low) and assessed the knowledge retained by various language models on this benchmark.
-Unlike previous studies, Sun et al.
-(2023) reported both accuracy (the percentage of questions answered correctly) and hallucination rate (the percentage of questions answered wrongly).
-They showed models are particularly bad at answering questions from medium- and low-frequency facts.
-Our work also examines whether LLMs can be used in lieu of more traditional KBs.
-Compared to previous research, we proposed a comprehensive evaluation framework which assesses not only the ability of LLMs to recall seen knowledge but also their ability to respond in the face of unseen knowledge.
-In addition, we evaluate LLM consistency when answering questions about identical knowledge.
-7 Conclusion
-In this paper, we defined a set of criteria that LLMs functioning as KBs should meet, focusing on factuality and consistency.
-We proposed various metrics operationalizing these criteria and used them to assess LLM performance when answering questions pertaining to both seen and unseen knowledge.
-We evaluated 26 popular LLMs and found that gpt-3.5-turbo is the most reliable among them.
-Additionally, we examined the impact of model size, fine-tuning, and ICL.
-Fine-tuning and ICL with unsure shots were shown to improve LLM capabilites when providing responses to questions about unseen knowledge, but do not significantly improve performance on seen knowledge.
-Increasing model size boosts performance on seen knowledge but at the expense of performance on unseen knowledge.
-Notably, neither ICL nor fine-tuning were successful at improving LLM consistency.
-This highlights the critical need for continued research to develop more robust strategies that ensure both factuality and consistency, enabling LLMs to effectively function as KBs.
-Appendix A Datasets
-A.1 SeenQA Dataset
-SeenQA is composed of questions selected from the following datasets:
-  1. 1.
-    Natural Questions Kwiatkowski et al.
-(2019): This dataset includes questions sourced from web queries, each paired with a corresponding Wikipedia article containing the answer.
-The paper on Natural Questions was submitted to TACL in April 2018.
-  2. 2.
-    TriviaQA Joshi et al.
-(2017): This dataset comprises questions from Quiz League websites, supplemented by web pages and Wikipedia searches that may contain the answer.
-The paper on TriviaQA was submitted to Arxiv in May 2017.
-For this project, we focus only on questions supported by Wikipedia.
-  3. 3.
-    PopQA Mallen et al.
-(2023): This dataset targets long-tail entities.
-The authors used the Wikipedia dump from December 2018 in the retrieval augmented baseline, indicating that the knowledge in PopQA can be covered by the Wikipedia dump from that date.
-Wikipedia is a common source in the pre-training data of large language models (LLMs).
-Comparing the knowledge cutoff dates provided in Table 4, we can deduce that the knowledge involved in these three datasets must have been seen during training by the LLMs used in our study.
-SeenQA contains 3000 questions selected through a three-step process:
-  1. 1.
-    Factoid Question Extraction: We filter out questions starting with "why," those with multiple different answers, or answers longer than five tokens.
-  2. 2.
-    Removal of Time-Sensitive Questions: We use gpt-4-1106-preview to detect questions whose answers may change over time, using the prompt shown in Table 5.
-Such questions are then removed from the dataset.
-  3. 3.
-    Sampling: Finally, we randomly sample 1,000 questions from each dataset’s test or development set, if the test set is unavailable, discarding the supporting context to adapt to a closed-book setting.
-A.2 UnseenQA Dataset
-UnseenQA was created using 20 templates, shown in Table 3, covering five answer types: number, person, time, location, and others.
-These templates wet=re hand-written by the authors.
-Templates T1, T2, T5, T6, and T7 are about future events, so their answers are unknown at the time of writing.
-The remaining templates involve fictional persons whose names and related information cannot be found on the Internet.
-The templates include three types of placeholders:
-  1. 1.
-    Country/Region: We collected 150 country/region names from the list of National Olympic Committees on a Wikipedia page.
-  2. 2.
-    Medal Event: We collected 150 medal events from the official programme of the Olympic games, Paris 2024.
-  3. 3.
-    Person: We randomly sampled 150 person names from combinations of 100 first names, 100 middle names, and 100 last names.
-We manually checked these names on Google to ensure they have no Internet presence.
-UnseenQA consists of 3000 questions, with 150 questions per template. This dataset was created on April 13, 2024. Therefore, LLMs trained on knowledge before this date should not know the answers.
+We evaluate 26 popular LLMs, including gpt-3.5-turbo, Flan-t5, llama1, llama2, llama3, mistral, gemma, and Phi2.
+Detailed descriptions of the evaluated LLMs are provided in Table 4 in Appendix B.
+Detailed descriptions are available in Table 4 in Appendix B.
+We test LLMs of different sizes: small (0.08B–3B), medium (7B–13B), and large (65B–70B).
+"Fine-tuned LLMs" refer to those fine-tuned via instruction-tuning or reinforcement learning from human feedback (e.g., llama3instruct-8B), while "base LLMs" refer to models without fine-tuning (e.g., llama3-8B).
 Answer Type Abb Template
@@
 Table 1: Question templates used to create UnseenQA
-Appendix B LLMs used in this project
+3.2 Datasets
+SeenQA
+SeenQA is a composite dataset comprising 3,000 questions sourced from the test sets (or development sets, where test sets were unavailable) of Natural Questions, TriviaQA, and PopQA (see Appendix A).
+All three datasets are derived from Wikipedia, with a knowledge cutoff date no later than December 2018.
+Given that Wikipedia is a frequent source for pre-training LLMs and the evaluated LLMs in this study have a knowledge cutoff date beyond April 2019 (as shown in Table 4), it can be inferred that the knowledge in these datasets is available to LLMs during their training.
+The creation of SeenQA involved a three-step process: 1) Factoid Question Extraction: exclude "why" questions, those with multiple answers, or answers exceeding five tokens.
+2) Time-Sensitive Question Removal: use gpt-4-1106-preview (prompt in Table 5) to detect and remove questions with time-variant answers.
+3) Random Sampling: randomly select 1,000 questions per source, discarding supporting context for a closed-book setting.
+UnseenQA
+UnseenQA is a new QA dataset designed to ensure the LLMs tested in this study lack access to knowledge required to answer questions.
+It consists of 3,000 questions generated from 20 hand-written templates (150 questions per template), as detailed in Table 1, spanning five answer types: number, person, time, location, and others.
+Templates T1–T7 focus on future events with unknown answers at the time of writing, while the remaining templates involve fictional individuals whose names and details are not available online.
+The templates feature three placeholder types: 1) Country/Region: 150 names sourced from the National Olympic Committees listed on the Wikipedia page.
+2) Medal Event: 150 medal events from the official programme of the Olympic games, Paris 2024.
+3) Person: 150 randomly generated names from combinations of 100 first, middle, and last names, manually verified to lack online presence.
+The dataset was created on April 13, 2024.
+All LLMs studied were released before this date and thus lacked access to the knowledge.
+3.3 Evaluation on a Single Response
+Uninformative
+We classify uninformative responses into three categories: repetition, none, and unsure (see Appendix C).
+Repetition refers to responses that repeatedly echo a specific string.
+We detect this using regular expressions and word frequency analysis.
+None includes responses that lack relevant content, such as empty strings or those merely repeating the question.
+Unsure applies to responses where the model says it is unable to answer or does not know.
+Examples include phrases like "I am not sure" or "I am just an AI", etc.
+Correct
+A response is considered correct if it is an exact match with the ground truth or is similar enough as judged by GPT-4o.
+Following prior work Sun et al.
+(2023), which reported a 98% agreement rate between ChatGPT and human evaluations, we adopt their evaluation prompt (detailed in Table 7 in Appendix D).
+Figure 2: Top 15 LLMs ranked by (a) Net Correct Rate, (b) Uninformative Rate, (c) $C_{\textit{correct}}$, (d) $C_{\textit{wrong}}$, (d) Net Consistent Correct Rate, (e) Inconsistent/Uninformative Rate.
+Values are scaled by 100 (full results in Appendix E).
+Consistency Score
+To compute $\text{Cons}(q,r)$, we set $X_{\text{MCQs}}$ (total number of MCQs) to 20, and each MCQ includes question $q$ and 5 options (the original response $r$, 3 random distractor options, and an ‘unsure’ option).
+3.4 Prompts and Hyper-parameters
+All LLMs were evaluated using greedy decoding (temperature 0 for gpt-3.5-turbo) with a maximum of 100 new tokens.
+To provide a comprehensive evaluation, we experimented with three types of prompt settings: zero-shot, four-shot, and four-shot with two unsure shots.
+To avoid any bias introduced by fixed examples, we employed a dynamic few-shot method following the work of Nori et al.
+(2023).
+For further details of the prompt settings, refer to Appendix D.
+4 Quantitative Results
+We present results for all LLMs under different prompt settings in in Appendix E (Table 9: factuality, Table 10: consistency, and Table 11: reliability).
+LLM rankings based on different metrics are shown in Figure 6 and Figure 7, and also in Appendix E.
+In the remainder, we discuss our results in dynamic four-shot settings.
+Which LLM is most factual?
+As shown in Figures 2 (a) and (b), gpt-3.5-turbo performs best on seen knowledge, while gemma-instruct-7B excels at unseen knowledge.
+gpt-3.5-turbo and llama2chat-70b emerge as the most balanced models, demonstrating top performance across both seen and unseen knowledge.
+Notably, Figure 2 (a) highlights a limitation of CR, the standard metric for factuality.
+For instance, while llama1-13b has a significantly higher CR than gemma-instruct-7B, their behavior differs in terms of wrong responses.
+llama1-13b exhibits a 33.50% higher WR rate, leading to a lower NCR compared to gemma-instruct-7B.
+Which LLM is most consistent?
+As shown in Figures 2 (c) and (d), llama3instruct-70b is most consistent on questions it answers correctly, while llama2chat-7b is least consistent on questions it answers incorrectly.
+Among these models, llama3instruct-8B is most balanced, with a high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$.
+Which LLM is most reliable?
+As shown in Figures 2 (e) and (f), gpt-3.5-turbo excels at seen knowledge, while gemma-instruct-7B performs best on unseen knowledge.
+Among these models, gpt-3.5-turbo demonstrates balanced performance, with a top NCCR and a high IUR.
+Note that a more factual LLM is not necessarily more reliable.
+For example, llama3-instruct-8B ranks below the 15th in NCR and 13th in NCCR due to its relatively high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$.
+5 Discussion
+5.1 Correlation Analysis
+Comparisons Zero-Shot Four-Shot Four-Two
+NCR vs. UR 0.27 0.34 0.62
+NCCR vs. IUR --0.17 --0.12 0.41
+$C_{\textit{correct}}$ vs. $C_{\textit{wrong}}$ 0.81 0.78 0.51
+NCR vs. $C_{\textit{correct}}$ 0.60 0.41 0.37
+NCR vs. $C_{\textit{wrong}}$ 0.48 0.64 0.41
+UR vs. $C_{\textit{correct}}$ 0.43 0.07 0.35
+UR vs. $C_{\textit{wrong}}$ 0.34 0.05 0.48
+Table 2: Pairwise correlation of LLM performance on different metrics under different prompt settings.
+The correlations are computed across all LLMs (26 data points).
+We report Pearson’s $\rho$, with underlined values indicating statistical significance ($p<0.05$).
+Four-Two refers to the four-shot setting with two unsure shots.
+Does an LLM with strong performance on seen knowledge tend to perform strong on unseen knowledge?
+We next examine whether there is a linear relationship between performance on seen and unseen knowledge.
+Performance on seen knowledge does not reliably predict performance on unseen knowledge in zero-shot and few-shot settings (without unsure shots).
+Table 2 (the first and second rows) shows the correlation values (Pearson’s $\rho$) between NCR and UR, and NCCR and IUR, revealing no statistically significant correlations in these settings.
+However, in the four-shot setting with two unsure shots, correlations across all metrics are significant.
+While seen knowledge performance does not transfer to unseen knowledge, specific prompt manipulations can improve these correlations (see the last column in Table 2).
+Does an LLM with high consistency score on correctly answered questions tend to perform less consistent on wrongly answered questions?
+As shown in Figure 11 and Table 10 in Appendix E, an LLM tends to be more consistent on questions it answers correctly than on those it answers wrongly.
+However, when comparing different LLMs, models with higher $C_{\textit{correct}}$ also tend to have higher $C_{\textit{wrong}}$, as indicated by the positive, significant correlation reported in Table 2 (the third row).
+This finding suggests that LLMs consistent with correct responses are also consistent with wrong responses, contradicting the expectation of high $C_{\textit{correct}}$ and low $C_{\textit{wrong}}$.
+This highlights a notable flaw in current models, which future work should address.
+Does strong factuality performance correlate with strong consistency performance?
+As shown in Table 2 (last four rows), there is a positive correlation between NCR/UR and $C_{\textit{correct}}$/$C_{\textit{wrong}}$, particularly in zero-shot settings.
+This suggests that LLMs with strong factuality performance tend to be confident in their responses, maintaining consistency whether those responses are correct or wrong.
+5.2 Model Size, Fine-tuning and ICL Impact
+Figure 3: The impact of (a) model size, (b) fine-tuning, (c) ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR.
+Different metrics are color-coded.
+See Appendix E for more detailed visualization.
+Values are scaled by 100.
+Figure 4: The impact of question type on LLM performance, measured by Uninformative Rate (UR) on unseen knowledge (scaled by 100).
+Questions are grouped by answer type.
+Higher values have darker shades.
+How does model size affect LLMs performance?
+As shown in Figure 3(a), larger LLMs perform better on seen knowledge but worse on unseen knowledge.
+As model size increases, both NCR (blue line) and NCCR (purple line) improve, indicating better performance on questions related to seen knowledge.
+However, while the UR (orange line) increases, the IUR (brown line) decreases.
+This suggests that larger models make fewer wrong responses on unseen knowledge but tend to give more consistent wrong responses.
+We also observe that larger LLMs are more consistent, even with wrong responses.
+Both $C_{\textit{correct}}$ (green line) and $C_{\textit{wrong}}$ (red line) rise significantly with model size.
+While higher consistency in correct responses is desirable, the increase in consistency for wrong responses presents a risk.
+Larger models may confidently and consistently produce incorrect yet convincing information, heightening the potential for misinformation if not carefully managed.
+How does fine-tuning affect LLMs performance?
+Figure 3(b) suggests that fine-tuning improves performance on unseen knowledge but degrades performance on seen knowledge.
+After fine-tuning, both UR (orange line) and IUR (brown line) increase significantly, indicating improved handling of unseen knowledge.
+However, the decline in NCR (blue line) and NCCR (purple line) shows that fine-tuning reduces the model’s effectiveness with seen knowledge.
+We also see that fine-tuning has no impact on consistency.
+There is no notable change in $C_{\textit{wrong}}$ (red line) or $C_{\textit{correct}}$ (green line) after fine-tuning which suggests that current instruction-tuning and RLHF methods do not improve LLM consistency.
+How does ICL affect LLMs performance?
+Figure 3(c) shows that ICL does not improve performance on seen knowledge, but unsure shots enhance performance on unseen knowledge.
+ICL does not increase NCR (blue line) or NCCR (purple line), which contrasts with previous findings on LLMs’ ICL ability Brown et al.
+(2020); Chada and Natarajan (2021); Touvron et al.
+(2023); Bai et al.
+(2023).
+Prior research largely focuses on the correct rate in few-shot settings without unsure shots.
+As shown in Table 9 and Figure 12 (Appendix E), while the four-shot setting significantly increases the correct rate, it also raises the wrong rate.
+Adding unsure shots helps reduce the wrong rate but also lowers the correct rate.
+For unseen knowledge, incorporating two unsure shots in the four-shot setting substantially improves UR (orange line) and IUR (brown line), indicating better handling of unknown questions.
+ICL with unsure shots reduces consistency in wrong responses.
+Adding two unsure shots in the four-shot setting decreases $C_{\textit{wrong}}$ (red line).
+Without unsure shots, there is no significant change in $C_{\textit{wrong}}$ (red line) or $C_{\textit{correct}}$ (green line).
+5.3 Behavior on Unseen Knowledge
+How do LLMs perform on different types of questions about unseen knowledge?
+Figure 4 shows that base LLMs tend to overestimate their knowledge when answering numerical and temporal questions.
+Their UR scores are significantly lower for these types of queries, indicating a tendency to provide misleading responses even when they lack relevant knowledge.
+6 Related Work
+Petroni et al.
+(2019) first explored using pre-trained LMs as KBs, introducing LAMA and showing that BERT retains relational knowledge with precision as the metric.
+Roberts et al.
+(2020) evaluated knowledge storage and retrieval using natural language queries, measuring accuracy.
+Wang et al.
+(2021) fine-tuned BART with related passages to instill factual knowledge, assessing masked span recovery accuracy.
+He et al.
+(2024) trained T5 and LLaMA2 on Wikidata to evaluate large-scale knowledge retention via exact match and F1 scores.
+Sun et al.
+(2023) tested LLMs on 18,000 fact-based QA pairs, reporting both accuracy and hallucination rates.
+Many studies have highlighted the issue of factuality in current LLMs.
+To address this, new benchmarks have been proposed to assess factuality Muhlgay et al.
+(2024); Zhao et al.
+(2024); Liu et al.
+(2024), although these still rely on correct rate as the primary metric.
+Several methods to enhance LLM factuality have also been introduced Wang et al.
+(2022); Hase et al.
+(2024); Cohen et al.
+(2024); Qin et al.
+(2024).
+However, evaluating these improved models falls outside the scope of our experiments and is suggested for future work.
+Research on consistency Rajan et al.
+(2024); Sreekar et al.
+(2024); Saxena et al.
+(2024) has shown that LLMs often struggle with providing consistent responses.
+These works, however, do not differentiate between the consistency expectations for correctly and wrongly answered questions.
+Compared to previous research, we propose a comprehensive framework to evaluate not only whether LLMs recall seen knowledge but also their ability to respond to unseen knowledge.
+In addition, we evaluate LLM consistency when answering questions about identical knowledge.
+7 Conclusion
+In this paper, we rethink the requirements for evaluating LLMs as KBs and propose criteria emphasizing factuality and consistency, and the combination thereof which we argue is an indicator of reliability.
+We proposed various metrics operationalizing these criteria and used them to assess LLM performance when answering questions pertaining to both seen and unseen knowledge.
+We evaluated 26 LLMs on our newly proposed SeenQA and UnseenQA datasets, and examined the impact of model size, fine-tuning, and ICL.
+Our experimental results highlight the critical need for continued research to develop more robust strategies that ensure both factuality and consistency, enabling LLMs to reliably function as KBs.
+Limitations
+First, due to budget constraints, we conducted in-depth evaluations on only a single closed-source LLM (GPT-3.5-turbo).
+Nevertheless, this limitation does not diminish the contributions of our work.
+Our study includes a broad comparative analysis of 26 different LLMs, ensuring that the insights gained are comprehensive and not confined to any single model.
+Furthermore, the primary objective of our paper is to introduce a systematic framework for evaluating LLM-as-KB.
+This framework is universally applicable, offering value for the evaluation of a wide range of LLMs beyond those specifically analyzed in this study.
+Second, our evaluation focuses primarily on factoid questions, which assess the models’ ability to recall specific factual knowledge rather than perform complex reasoning.
+Investigating how LLMs handle such complex queries remains an important direction for future research.
+Appendix A Datasets
+SeenQA is composed of questions selected from the following open-sourced datasets:
+  1. 1.
+    Natural Questions Kwiatkowski et al.
+(2019): This dataset includes questions sourced from web queries, each paired with a corresponding Wikipedia article containing the answer.
+The paper on Natural Questions was submitted to TACL in April 2018.
+  2. 2.
+    TriviaQA Joshi et al.
+(2017): This dataset comprises questions from Quiz League websites, supplemented by web pages and Wikipedia searches that may contain the answer.
+The paper on TriviaQA was submitted to Arxiv in May 2017.
+For this project, we focus only on questions supported by Wikipedia.
+  3. 3.
+    PopQA Mallen et al.
+(2023): This dataset targets long-tail entities.
+The authors used the Wikipedia dump from December 2018 in the retrieval augmented baseline, indicating that the knowledge in PopQA can be covered by the Wikipedia dump from that date.
+Wikipedia is a common source in the pre-training data of large language models (LLMs).
+Comparing the knowledge cutoff dates provided in Table 4, we can deduce that the knowledge involved in these three datasets must have been seen during training by the LLMs used in our study.
+Uninformative Type Responses Examples
+repetition
+1. (a) (b) (c) (d) (e) (f) (g) (h) (i) (j) (k) (l) (m) (n) (o) (p) (q) (r) (s) (t) (u) (v) (w) (x) (y) (z)
+(aa) (ab) (ac) (ad) (ae) (af) (
+1. 2. 3. 4. 5. 6. 7. 8. 9. 10. 11. 12. 13. 14. 15. 16. 17. 18. 19. 20. 21. 22. 23. 24. 25. 26. 27.
+a swollen swollen swollen swollen swollen swollen swollen swollen swollen swollen
+swollen swollen swollen swollen swollen swollen swollen swollen
+none
+noah catherine cooper’s mom is ________.
+<a href="https://i.stack.imgur.com/88888.png" rel="nofollow noreferrer"><image></a>
+unsure
+unsure
+i do not know.
+i’m just an ai, i don’t have access to real-time information or the ability to predict the future
+Table 3: Uninformative Responses Examples
+Appendix B LLMs Used
 Table 4 summarizes the LLMs used in our experiments.
@@
 Llama 1’s pre-training data includes Wikipedia dumps from June to August 2022.
-Appendix C Prompts used in our Experiments
+Appendix C Uninformative Responses Examples
+Table 3 provides some examples of uninformative responses.
+As illustrated in Figure 5 (fourth column), fine-tuned LLMs are able to explicitly acknowledge their lack of knowledge by answering ‘unsure’ to questions about unseen knowledge.
+In contrast, base LLMs often produce responses classified as ‘none’ or ‘repetition’ in the absence of unsure shots (contrast columns one and two with column three in Figure 5).
+Figure 5: Distribution of uninformative responses given by LLMs to questions about unseen knowledge.
+We report results for the llama3-8b, gemma-7b, and their fine-tuned models (fourth column) but observe similar trends on other models (omitted for the sake of brevity).
+Appendix D Prompts Used
+To provide a comprehensive evaluation, we experimented with three types of prompt settings: zero-shot, four-shot, and four-shot with two unsure shots.
+To avoid any bias introduced by fixed examples, we employed a dynamic few-shot method following the work of Nori et al.
+(2023).
+We collected two repositories, $R_{seen}$ and $R_{unseen}$.
+$R_{seen}$ includes 280 question-answer pairs about seen knowledge (200 from the unused data of PopQA and training data of Natural Questions and TriviaQA; 80 are generated using the templates in Table 1).
+$R_{unseen}$ consists of 40 question-answer pairs about unseen knowledge, all generated using the templates in Table 1.
+We used text-embedding-3-small to embed the questions in the repositories and test questions as vector representations.
+For each test question under the four-shot setting, we retrieved its nearest four questions from $R_{seen}$.
+Under the four-shot with two unsure shots setting, we retrieved the nearest two questions from $R_{seen}$ and two from $R_{unseen}$.
 The prompt used to detect time-sensitive questions is shown in Table 5.
@@
 Table 8: The prompt used to generate distractors for consistency tests.
-Appendix D Full Experimental Results
+Appendix E Full Experimental Results
 Table 9, Table 10, and Table 11 provides the detailed results of LLMs’ performance on factuality, consitency, and reliablity respectively.
@@
 Figure 7 compare different LLMs’ factuality, consistency, and reliability performance.
+Figure 8,  9, and  10 provide detailed illustration of the impact of model size, fine-tuning, and ICL.
+Figure 11 shows $C_{\textit{correct}}$ and $C_{\textit{wrongs}}$ performance of LLMs ranked in $C_{\textit{correct}}$.
+Figure 12 shows correct rate and wrong ate under different prompt settings for LLMs ranked in correct rate under zero-shot.
 Model Params zero-shot four-shot four-shot-2
@@
 Phi2 3B 65.97 21.43 -44.53 13.83 72.10 21.40 -50.70 14.77 62.07 19.33 -42.73 50.20
-Table 9: Factuality performance.
+Table 9: Factuality performance (Values are scaled by 100).
 Model Params zero-shot four-shot four-shot-2
@@
 Phi2 3B 30.09 18.27 24.18 54.92 37.21 19.02 28.11 67.34 38.48 15.78 27.13 68.16
-Table 10: Consistency performance.
+Table 10: Consistency performance (Values are scaled by 100).
 Model Params zero-shot four-shot four-shot-2
@@
 Phi2 3B 19.85 11.77 -8.08 84.25 26.83 14.41 -12.42 83.79 23.88 13.17 -10.71 92.14
-Table 11: Reliability performance.
+Table 11: Reliability performance (Values are scaled by 100).
 Figure 6: Ranking of LLMs based on different Metrics.
 Figure 7: Visualization of LLMs’ factuality, consistency and reliablity performance.
+Figure 8: The impact of model size on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR (values are scaled by 100).
+Different metrics are color-coded.
+LLMs are shown in three sizes, small, medium, and large and are grouped into ‘base’ and fine-tuned ones.
+Figure 9: The impact of fine-tuning on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR (values are scaled by 100).
+Different metrics color-coded.
+This analysis only considers the performance of Llama2, Llama3, Mistral, and Gemma as these families include both base LLMs and fine-tuned versions.
+Models are shown in three sizes, small, medium, and large.
+Figure 10: The impact of ICL on LLM performance, measured with NCR, UR, Ccorrect, Cwrong, NCCR, and IUR (values are scaled by 100).
+Different metrics are color-coded.
+We compare zero-shot and four-shot settings; and zero-shot against four-shot with two unsure shots.
+LLMs (‘base’ and fine-tuned ones) are in three sizes, small, medium, and large.
+Figure 11: $C_{\textit{correct}}$ and $C_{\textit{wrongs}}$ performance of LLMs ranked in $C_{\textit{correct}}$.
+Figure 12: Correct Rate and Wrong Rate under different prompt settings for LLMs ranked in Correct Rate under zero-shot.